### v1

# Deep Transfer Learning: Model Framework and Error Analysis

Abstract
This paper presents a framework for deep transfer learning, which aims to leverage information from multi-domain upstream data with a large number of samples ($n$) to a single-domain downstream task with a considerably smaller number of samples ($m$), where $m\ll n$, in order to enhance performance on downstream task. Our framework has several intriguing features. First, it allows the existence of both shared and specific features among multi-domain data and provides a framework for automatic identification, achieving precise transfer and utilization of information. Second, our model framework explicitly indicates the upstream features that contribute to downstream tasks, establishing a relationship between upstream domains and downstream tasks, thereby enhancing interpretability. Error analysis demonstrates that the transfer under our framework can significantly improve the convergence rate for learning Lipschitz functions in downstream supervised tasks, reducing it from $\tilde{O}(m^{-\frac{1}{2(d+2)}}+n^{-\frac{1}{2(d+2)}})$ (“no transfer”) to $\tilde{O}(m^{-\frac{1}{2(d^{*}+3)}}+n^{-\frac{1}{2(d+2)}})$ (“partial transfer”), and even to $\tilde{O}(m^{-1/2}+n^{-\frac{1}{2(d+2)}})$ (“complete transfer”), where $d^{*}\ll d$ and $d$ is the dimension of the observed data. Our theoretical findings are substantiated by empirical experiments conducted on image classification datasets, along with a regression dataset.
1 Introduction
Training a model solely on one or multiple analogous tasks can impede its performance when applied to a distinct yet similar task [1, 2], due to the presence of data bias and task bias. By recognizing that multiple diverse tasks may share a common representation, transfer learning emerges as a crucial solution. The relevant literature on transfer learning [3, 4] provides valuable insights into its applications and discussions. However, the existing literature lacks systematic research on the following issues, particularly within the framework of deep learning:
  • (i)
    How to learn a sufficient and invariant representation form the upstream data?
  • (ii)
    Given that the representations of different datasets should not be identical and there should be shared as well as specific features, how can we design a model framework to accurately characterize this?
  • (iii)
    When the upstream representation is sufficient, which features are relevant for downstream prediction? And how should transfer learning be described when the upstream representation is not sufficient for the downstream task?
In this paper, we aim to address the aforementioned questions by proposing a deep learning framework for transfer learning. The goal is to derive feature representations from multi-domain upstream data, characterized by a large sample size, and apply them to a single-domain downstream task with a considerably smaller number of samples, in order to improve performance on the downstream task. Our model accomplishes this in three key steps. Firstly, it permits the presence of common and distinct features across domains in the upstream data and automatically identifies them. Secondly, it explicitly expresses which upstream features contribute to the downstream tasks, establishing connection among domains and enhancing interpretability. Moreover, a $d$-dimensional nonlinear function $Q^{*}$ is introduced for the downstream task to learn features that are independent of the upstream representation, determining the level of difficulty in transfer learning. Particularly, in the case where $Q^{*}=0$, the upstream representation is sufficient to the downstream task, transfer from upstream to downstream is facile, leading to a parametric convergence rate. On the other hand, when specific information is encoded in the non-zero $Q^{*}$, the transfer problem becomes more challenging. Nevertheless, with the aid of the representation learned from the upstream data, our task is simplified to the estimation of an independent $Q^{*}$ of upstream representation, resulting in a convergence rate that surpasses the pace of the Empirical Risk Minimization (ERM) estimator for the downstream task. Importantly, we emphasize that our model framework seamlessly adapts to these scenarios, accommodating varying degrees of information transfer and yielding distinct convergence rates for the downstream task.
The main contributions in this paper are as follows:
  • We introduce an adaptive deep transfer learning framework. Initially, we acquire deep representations from the upstream dataset through identifiable models, ensuring sufficiency and invariance. In instances of “complete transfer”, where the acquired representations offer sufficient information for downstream prediction, we only need to learn an $r$-dimensional vector function for the downstream task. In case that “partial transfer”, an additional learning for a low dimensional function is required.
  • Our error analysis reveals that the convergence speed, when learning a $\beta$-smooth function with over-parameterized neural networks for the downstream task, can manifest at distinct rates: slow $\mathcal{O}(m^{-\frac{\beta}{2(d+1+2\beta)}})$, moderate $\mathcal{O}(m^{-\frac{\beta}{2(d^{*}+1+2\beta)}})$ with $d^{*}\ll d$ or fast $\mathcal{O}(m^{-1/2})$. These rates correspond to “no transfer” that the downstream task necessitates learning a $d$-dimensional function, “partial transfer” problems where the downstream task entails learning an $d^{*}$-dimensional function, and “complete transfer” that the downstream task involves solving a sparse-penalized least squares problem, respectively.
1.1 Related Work
Extensive research has been conducted on single-source transfer learning under covariate shift and posterior shift, encompassing both parametric and nonparametric models. Theoretical investigations into transfer learning within parametric models, such as high-dimensional linear regression [5] or generalized linear models [6], have been thoroughly explored. Nonparametric transfer learning in regression with posterior shift has been addressed by [7, 8], as well as in the work of [9]. This area has also been explored by [10, 11] in the scenario of covariate shift. Nonparametric transfer learning in classification, has been studied in general by [12, 13, 14]. Theoretical results and adaptive procedures have been developed for both the covariate shift setting [15, 16] and the posterior drift setting [17, 18]. Minimax rates are derived under posterior drift by [8] and under covariate shift by [19], respectively. However, it is noteworthy that even in the extensive theoretical literature mentioned above, deep transfer learning, even for the single-source (a special case of our study with $p=1$), has not been thoroughly explored. Non-parametric analysis of transfer learning involving shared representations across diverse tasks have been explored, see [20, 21, 22, 23] and the reference therein. In downstream task, task-specific linear functions and nonlinear functions have been studied in [20, 22] and [21, 23], respectively. In these studies, their theories consistently demonstrate the improvement of transfer learning in terms of sampling efficiency under the assumption that sufficient representations of downstream task can be obtained from the upstream domains. In addition, these above mentioned theory is developed for well-specified neural network models without considering the approximation error as we did in this paper.
The concept of a shared invariant representation across domains is not only applicable in transfer learning but also extends to various related domains and tasks, including domain adaptation [24, 25, 26], domain generalization [27, 28, 29, 30], and meta-learning [31]. Recently, [32] provided theoretical insights into domain-invariant representations within the context of multiple-source domain adaptation and domain generalization. In this paper, we learned deep representations from the upstream dataset through identifiable models, ensuring not only invariance but also sufficiency. Another point sets our method apart from the aforementioned ones is that we allow for the existence of not only shared representations but also specific features, making our model more flexible and practical. Additionally, the loss function we used to train this representation differs significantly from the above ones.
Learning a pre-trained model with huge data set and then fine-tuning on specific task [33, 34, 35, 36, 37, 38, 39, 40, 41, 42] has empirically proven to be a highly effective strategy in recent years. Our framework for multi-domain transfer learning can also be interpreted along these lines, wherein learning representations from multiple domains can be viewed as pre-training, and refining downstream predictions can be seen as fine-tuning. Consequently, the theoretical results established in this paper also contribute to enhancing our understanding of pre-training and fine-tuning processes.
The remainder of this paper is structured as follows: Preliminary details about the framework, including essential definitions and model setup, are provided in Section 1. Our main theoretical results are presented in Section 3. Experimental results are showcased in Section 4, followed by a summary in Section 5. Technical proofs are provided in the appendix.
2 Preliminaries
Table 1: Summary of notations
$\mathcal{D}=\{(\mathbf{X}_{i},Y_{i},S_{i})\}_{i=1}^{n}$ dataset for upstream
$\mathbb{P}(\mathbf{x},y,\mathbf{s})$ the unknown distribution of upstream data
$(\mathbf{X},Y,S)$ r.v.s (random variables) for upstream data
$\mathcal{X}$,$\mathcal{Y}$,$\mathcal{S}$ support of upstream data $\mathbf{X}$, $Y$, $\mathbf{s}$ respectively
$\mathcal{T}=\{(\mathbf{X}_{T,i},Y_{T,i})\}_{i=1}^{m}$ dataset for downstream
$\mathbb{P}_{0}$ the unknown distribution of downstream data
($\mathbf{X}_{T},Y_{T}$) the pair of r.v.s (random variables) $(X_{T},Y_{T})$ with its joint distribution $\mathbb{P}_{0}$
$\varepsilon_{1},\varepsilon_{2}$ bounded random noises: $|\varepsilon_{i}|\leq 1,i\in[2]$
$\mathcal{U}_{r}$ a random vector with each coordinate being mutually independent and
have uniform distribution $U[0,1]$
$\{\xi_{i}\}_{i=1}^{n}$ the i.i.d random samples drawn from $\mathcal{U}_{r}$
$n$ the number of upstream data
$m$ the number of downstream data
$p$ the number of different domain
$n_{s}$ the number of upstream data from domain $s$
$d$ the dimension of input data $\mathbf{X}$
$d^{*}$ the reduced dimension of downstream input data $\mathbf{X}_{T}$
$r$ the dimension of intrinsic representation
($\mathbf{X},Y$) the combine r.v.s(random variables) pairs with
[…]
$\mathbf{X}_{S}$, $Y_{S}$ the r.v.s(random variables) from domain $S$
$\mathbf{X}\perp\!\!\!\perp Y$ for r.v.s (random variables) $\mathbf{X}$ and $Y$, $\mathbf{X}$ is independent with $Y$
$\mathbf{X}\perp\!\!\!\perp Y\mid Z$ for r.v.s (random variables) $\mathbf{X}$, $Y$ and $Z$, condition on $Z$,
$\mathbf{X}$ is independent with $Y$
$\mathbf{h}^{*}_{s}(\mathbf{X})$ the oracle representation for domain $s$ with $\mathbf{X}_{s}\perp\!\!\!\perp Y_{s}\mid\mathbf{h}^{*}_{s}(\mathbf{X})$
$\mathbf{h}^{*}(\mathbf{X})$ the oracle representation satisfied $\mathbf{X}\perp\!\!\!\perp Y\mid\mathbf{h}^{*}(\mathbf{X})$ and $S\perp\!\!\!\perp\mathbf{h}^{*}(\mathbf{X})$
$\mathbf{F}^{*}_{S}$ the oracle sparse transformation vector for domain $S$
$\mathbf{F}_{T}^{*}$ the oracle sparse transformation vector for downstream data
$Q^{*}$ an additional $\mathbb{R}^{d}\to\mathbb{R}$ function to be learned at downstream task
and satisfy $Q^{*}(X_{T})\perp\!\!\!\perp\mathbf{h}^{*}(X_{T})$
$\mathbf{A}^{*}$ the oracle dimensionality reduction matrix for downstream input data
to control downstream task complexity
$q^{*}$ the oracle $\mathbb{R}^{d^{*}}\mapsto\mathbb{R}$ function satisfy $Q^{*}(X_{T})=q^{*}(\mathbf{A}^{*}X_{T})$
$\mathbf{F}\in\mathbb{R}^{pr}$ the shorthand for $(\mathbf{F}_{1},\ldots,\mathbf{F}_{p})$
[…]
$\|\cdot\|_{1}$ a vector’s $\ell_{1}$ norm or the sum of the $\ell_{1}$ norms of the columns of a matrix
$[q]$ shorthand for integer sets $\{1,\ldots,q\}$
$\|f\|_{L}$ the Lipschitz constant of function $f$
$\mathcal{H}$ a neural network(NN) function class of features mapping $\mathbb{R}^{d}\to\mathbb{R}^{r}$
$\mathcal{G}_{1}$,$\mathcal{H}_{1}$ a NN function class of lipschitz-constrained mapping $\mathbb{R}^{r}\to\mathbb{R}$
$\mathcal{Q}$ a NN function class during tuning mapping $\mathbb{R}^{d}\to\mathbb{R}$
$\lambda,\tau,\mu,\kappa,\chi,\zeta$ regularization parameters
$\mathcal{V}[\mathbf{h}(\mathbf{X}),S]$ a measure used to evaluate independence between $\mathbf{h}(\mathbf{X})$ and $S$,
in this work we use the distance covariance
$\widehat{\mathcal{V}}_{n}[\mathbf{h}(\mathbf{X}),S]$ the empirical form of $\mathcal{V}[\mathbf{h}(\mathbf{X}),S]$
$\mathcal{W}_{1}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})$ the 1-Wasserstein distance used to match the distribution of $\mathbf{h}(\mathbf{X})$ and
$\mathcal{U}_{r}$, in this work we use its dual form to calculate it
$\hat{\mathcal{W}_{1}}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})$ the empirical dual form of $\mathcal{W}_{1}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})$ by replacing the
searching function class by norm-constrained NN
$R(\mathbf{F},\mathbf{h})$ the population risk of learned function $(\mathbf{F},\mathbf{h})$;
i.e, $\mathbb{E}[\|Y_{S}-\mathbf{F}_{S}\mathbf{h}(\mathbf{X}_{S})\|^{2}]+\lambda\mathcal{V}[\mathbf{h}(\mathbf{X}),S]+\tau\mathcal{W}_{1}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})+\mu\|\mathbf{F}\|_{1}$
$\hat{R}(\mathbf{F},\mathbf{h})$ the empirical version of $R(\mathbf{F},\mathbf{h})$;
i.e, $\sum_{i=1}^{p}\sum_{j=1}^{n_{p}}\frac{1}{n}(Y_{i,j}-\mathbf{F}_{i}\mathbf{h}(X_{i,j}))^{2}+\lambda\widehat{\mathcal{V}}_{n}[\mathbf{h}(\mathbf{X}),S]+$
$\tau\hat{\mathcal{W}_{1}}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})+\mu\|\mathbf{F}\|_{1}$
$(\hat{\mathbf{F}},\hat{\mathbf{h}})$ the estimated sparse vector $\hat{\mathbf{F}}=(\hat{\mathbf{F}}_{1},\ldots,\hat{\mathbf{F}}_{p})$ and estimated feature
representation $\hat{\mathbf{h}}$; i.e, $(\hat{\mathbf{F}},\hat{\mathbf{h}})=\mathop{\rm argmin}_{(\mathbf{F},\mathbf{h})\in\mathcal{B}_{pr}(B)\circ\mathcal{H}}\hat{R}(\mathbf{F},\mathbf{h})$
$\ell(\cdot,\cdot)$ the general loss for regression or classification
[…]
^$\hat{\mathcal{L}}(\hat{\mathbf{h}},Q_{q,\mathbf{A}},\mathbf{F}_{T})$ the empirical risk of learned $(\hat{\mathbf{h}},Q_{q,\mathbf{A}},\mathbf{F}_{T})$;
i.e, $\sum_{i=1}^{m}\ell(\mathbf{F}_{T}\hat{\mathbf{h}}(\mathbf{X}_{T,i})+q(\mathbf{A}\mathbf{X}_{T,i}),Y_{T,i})/m+\kappa\hat{\mathcal{V}}_{m}[\hat{\mathbf{h}},Q_{q,\mathbf{A}}]+$
   $\chi\|\mathbf{F}_{T}\|_{1}+{\zeta\|\mathbf{A}\|_{F}^{2}}$
$\mathcal{L}(\mathbf{h},Q_{q,\mathbf{A}},\mathbf{F}_{T})$ the population risk of learned $(\hat{\mathbf{h}},Q_{q,\mathbf{A}},\mathbf{F}_{T})$;
i.e, $\mathcal{L}(\mathbf{h},Q_{q,\mathbf{A}},\mathbf{F}_{T})=\mathbb{E}_{X_{T},Y_{T}}[\hat{\mathcal{L}}(\mathbf{h}(X_{T}),q(\mathbf{A}X_{T}),\mathbf{F}_{T})]$
$\tilde{O}(\cdot)$ Asymptotic notation used to hide polylogarithmic factors in all relevant
parameters of the problem
2.1 Transfer learning based on sufficient and domain-invariant representation
2.1.1 Setup
[…]
$\mathbf{X}\perp\!\!\!\perp Y\mid\mathbf{h}^{*}(\mathbf{X}),$ (1)
$S\perp\!\!\!\perp\mathbf{h}^{*}(\mathbf{X}),$ (2)
where $\mathbf{X}\perp\!\!\!\perp Y\mid\mathbf{h}^{*}(\mathbf{X})$ indicates that the information of $\mathbf{X}$ to predict $Y$ is sufficiently encoded in $\mathbf{h}^{*}(\mathbf{X})$, and $S\perp\!\!\!\perp\mathbf{h}^{*}(\mathbf{X})$ signifies the need for the representation to be invariant across all domains in the upstream dataset. Under (1), we use the following regression model for identifying and estimating $\mathbf{h}^{*}(\mathbf{X})$:
$Y_{S}=\mathbf{F}^{*}_{S}\mathbf{h}^{*}(\mathbf{X}_{S})+\varepsilon_{1},\quad\mathbb{E}[\varepsilon_{1}]=0,$ (3)
where the sparse vector $\mathbf{F}^{*}_{S}\in\mathbb{R}^{1\times r}$ represents the transformation applied to the representation $\mathbf{h}^{*}(\mathbf{X}_{S})$ for domain $S$, $\varepsilon_{1}$ is a bounded noise with mean zero. The sparsity of $\mathbf{F}^{*}_{S}$ is used to identify the feature subset $\mathbf{h}^{*}_{S}(\mathbf{X}_{S})(\subset\mathbf{h}^{*}(\mathbf{X}_{S}))$ that capture the relationship between $\mathbf{X}_{S}$ and $Y_{S}$. The sparsity sets our framework apart from existing transfer learning methods by allowing features to vary or be shared across domains, rather than exclusively being shared. With the sparsity of $\mathbf{F}_{S}$, model (3) also provides a more parsimonious and interpretable representation for each domain. Domains with different feature sets enable us to better explore their interdomain differences. For the downstream data, we focus on single-target domain transfer learning. Let the data set $\mathcal{T}=\{(\mathbf{X}_{T,i},Y_{T,i})\}_{i=1}^{m}$ consist of $m$ samples of the random variable pair $(X{T},Y_{T})$, which are i.i.d. from an unknown distribution $\mathbb{P}_{0}$, which is also supported over $\mathcal{X}\times\mathcal{Y}$. We assume the supervised downstream task data following the regression model:
$Y_{T}=\mathbf{F}_{T}^{*}\mathbf{h}^{*}(X_{T})+Q^{*}(X_{T})+\varepsilon_{2},\quad\mathbb{E}[\varepsilon_{2}]=0,$ (4)
or the binary classification model
$Y_{T}\sim\mathrm{Ber}(\sigma(\mathbf{F}_{T}^{*}\mathbf{h}^{*}(X_{T})+Q^{*}(X_{T})))$ (5)
where $\varepsilon_{2}$ is bounded noise with zero mean, $\mathbf{F}_{T}^{*}$ is the sparse transformation vector for target domain and $Q^{*}:\mathbb{R}^{d}\to\mathbb{R}$ is an additional function to be learned which indicates that downstream tasks may not share sufficient similarity to predict $Y_{T}$ solely based on the representation $\mathbf{h}^{*}({X_{T}})$. Thus, $Q^{*}\neq 0$ serves the purpose of extracting additional information different from $\mathbf{h}^{*}(X_{T})$ which is implemented by forcing
$Q^{*}(X_{T})\perp\!\!\!\perp\mathbf{h}^{*}(X_{T}).$ (6)
To explicitly express the constraint (6), we further model
$Q^{*}(X_{T})=q^{*}(\mathbf{A}^{*}X_{T}),$ (7)
where $\mathbf{A}^{*}\in\mathbb{R}^{d}\rightarrow\mathbb{R}^{d^{*}}$ with $d^{*}\leq d$, and $q^{*}:\mathbb{R}^{d^{*}}\mapsto\mathbb{R}$. In fact, if $d^{*}=d$, setting $\mathbf{A}^{*}=\mathbf{I}$ reduces $q^{*}(\mathbf{A}^{*}X_{T})$ to $q^{*}(X_{T})$ or $Q^{*}(X_{T})$, indicating no restriction on the model (7). Therefore, the value of $d^{*}$ indicates the strength of the constraint (6). A smaller $d^{*}$ implies a more stringent constraint of (6). In practice, we can determine $d^{*}$ among various candidate integers by evaluating prediction accuracy.
When $\mathbf{F}_{T}^{*}=0$, the contribution of the upstream representation to the downstream task is negligible, and we refer to this scenario as “no transfer”. On the other hand, when $Q^{*}=0$, the entirety of upstream information can be seamlessly transferred to the downstream task. In this case of “complete transfer”, the task involves solely the estimation of the sparse vector $\mathbf{F}_{T}^{*}$ through penalized least squares. In situations where $\mathbf{F}^{*}_{T}\neq 0$ and $Q^{*}\neq 0$, we classify it as “partial transfer”. Hence, our downstream task model (4)-(7) can adaptively accommodate these three distinct scenarios. Furthermore, our Theorem 3, Remark 6 and Theorem 4 demonstrate that the end-to-end convergence rate distinguishes these settings.
Remark 1.
Practitioners have also explored the following transfer models,
$Y_{T}=f^{*}_{T}(\mathbf{h}^{*}(X_{T}))+\varepsilon_{2},\quad\mathbb{E}[\varepsilon_{2}]=0,$ (8)
or the binary classification model
$Y_{T}\sim\mathrm{Ber}(\sigma(f^{*}_{T}(\mathbf{h}^{*}(X_{T}))))$ (9)
where $f^{*}_{T}:\mathbb{R}^{r}\to\mathbb{R}$ is a low dimensional function to be learned nonparametrically in the downstream task. The model (8) or (9) suggests sufficient features from the upstream domains to the downstream task, as $f^{*}_{T}(\mathbf{h}^{*}(X_{T}))$ depends on $\mathbf{h}^{*}(X_{T})$. Here in (4) or (5) , we choose a linear sparse function over the nonparametric function $f^{*}_{T}(\cdot)$ for interpretability, as well as the consideration that the inherent nonparametric features are already captured by $\mathbf{h}^{*}(\mathbf{X}_{T})$.
2.1.2 Loss for representation learning
Notably, for any bijective mapping $\psi$ acting on $\mathbf{h}^{*}$, the composition $\psi\circ\mathbf{h}^{*}$ preserves the sufficiency and invariance properties required in (1)-(2). Therefore, we can constrain the distribution of $\mathbf{h}^{*}(\mathbf{X})$ to match the uniform distribution $\mathcal{U}_{r}$ without loss of generality, as the existence of such $\mathbf{h}^{*}(\mathbf{X})$ can be guaranteed by optimal transport theory [43]. Given that the entries of $\mathcal{U}_{r}$ are independent, we promote the disentanglement of $\mathbf{h}^{*}(\mathbf{X})$.
As discussed in Section 2.1.1, for different domain from upstream, only a portion of information from $\mathbf{h}^{*}(X_{S})$ is utilized during the prediction process. Consequently, we incorporate domain-specific extractors represented by the sparse vector $\mathbf{F}_{S}$, and add $\ell_{1}$ penalty in loss for the sparsity of $\mathbf{F}_{S}$. For simplicity, we denote $(\mathbf{F}_{1},\ldots,\mathbf{F}_{p})$ as $\mathbf{F}\in\mathbb{R}^{pr}$. Then, defining the population risk as:
$R(\mathbf{F},\mathbf{h})=$ $\mathbb{E}[\|Y_{S}-\mathbf{F}_{S}\mathbf{h}(\mathbf{X}_{S})\|^{2}]+\lambda\mathcal{V}[\mathbf{h}(\mathbf{X}),S]$ (10)
$+\tau\mathcal{W}_{1}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})+\mu\|\mathbf{F}\|_{1},$
where the first term is the regression loss for (3), and $\mathcal{V}[\mathbf{h}(\mathbf{X}),S]$ is a measure used to evaluate independence between $\mathbf{h}(\mathbf{X})$ and $S$ as required in (2), such as the distance covariance (detail definition see Section 2.2.1, and
$\mathcal{W}_{1}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})=\sup_{\|g\|_{L}\leq 1}\mathbb{E}_{\mathbf{x}\sim\mathbf{h}(\mathbf{X})}[g(\mathbf{x})]-\mathbb{E}_{\mathbf{y}\sim\mathcal{U}_{r}}[g(\mathbf{y})],$
is the Wasserstein distance used to match the distribution of $\mathbf{h}(\mathbf{X})$ and $\mathcal{U}_{r}$, and $\lambda$, $\tau,\mu$ are regularization parameters. Given data $\mathcal{D}=\{(\mathbf{X}_{i},Y_{i},S_{i})\}_{i=1}^{n}$, the empirical version of (10) can be written as:
$\hat{R}(\mathbf{F},$ $\mathbf{h})=\sum_{i=1}^{p}\sum_{j=1}^{n_{p}}\frac{1}{n}(Y_{i,j}-\mathbf{F}_{i}\mathbf{h}(X_{i,j}))^{2}+$ (11)
$\lambda\widehat{\mathcal{V}}_{n}[\mathbf{h}(\mathbf{X}),S]+\tau\hat{\mathcal{W}_{1}}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})+\mu\|\mathbf{F}\|_{1},$
where $\sum_{j=1}^{p}n_{p}=n$, and $\widehat{\mathcal{V}}_{n}[\mathbf{h}(\mathbf{X}),S]$ is the empirical form of $\mathcal{V}[\mathbf{h}(\mathbf{X}),S]$ (see Section 2.2.1 for detail), and
$\hat{\mathcal{W}_{1}}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})=\frac{1}{K_{3}}\sup_{g\in\mathcal{G}}\frac{1}{n}\sum_{i=1}^{n}(g(\mathbf{h}(\mathbf{X}_{i}))-g(\mathbf{\xi}_{i})),$
denotes the empirical form of the Wasserstein distance, wherein $\{\xi_{i}\}_{i=1}^{n}$ denotes i.i.d random samples drawn from $\mathcal{U}_{r}$ and the function class $\mathcal{G}$ is chosen as the norm-constrained neural network class whose Lipschitz constants are bounded by $K_{3}$, see Section 2.2.2 for detail.
We define the estimated sparse vector $\hat{\mathbf{F}}=(\hat{\mathbf{F}}_{1},\ldots,\hat{\mathbf{F}}_{p})$ and estimated feature representations $\hat{\mathbf{h}}$ as the following
$(\hat{\mathbf{F}},\hat{\mathbf{h}})=\mathop{\rm argmin}_{(\mathbf{F},\mathbf{h})\in\mathcal{B}_{pr}(B)\circ\mathcal{H}}\hat{R}(\mathbf{F},\mathbf{h}),$ (12)
where $\mathcal{B}_{pr}(B)$ is the ball in $\mathbb{R}^{pr}$ with radius $B$, and $\mathcal{H}$ is a deep neural network class with prescribed network structure. We summarize the procedure for representation learning from upstream data in Algorithm 1.
Algorithm 1 Upstream training
  Input: Upstream data $\{\mathbf{X}_{i},Y_{i},S_{i}\}_{i=1}^{n}$
  Initialize $\mathbf{F}$ and $\mathbf{h}$
  repeat
     $g\leftarrow\mathop{\rm argmax}_{g\in\mathcal{G}}\hat{\mathcal{W}_{1}}(\mathbf{h}(X),\mathcal{U}_{r})$
     $(\mathbf{F},\mathbf{h})\leftarrow\mathop{\rm argmin}_{(\mathbf{F},\mathbf{h})\in\mathcal{B}_{q}(B)\circ\mathcal{H}}\hat{R}(\mathbf{F},\mathbf{h})$
  until convergence
  Output: $\hat{\mathbf{F}},\hat{\mathbf{h}}$
2.1.3 Loss for downstream task
For the downstream learning, we plug in the learned $\hat{\mathbf{h}}$ in Algorithm 1 and fit the regression and classification models (4) (5) with downstream data $\mathcal{T}=\{(\mathbf{X}_{T,i},Y_{T,i})\}_{i=1}^{m}$, respectively. Specifically, we estimate the optimal $\mathbf{F}_{T}^{*}$ and $Q^{*}(\cdot)=q^{*}(\mathbf{A}^{*}\cdot)$ via
$(\hat{\mathbf{F}}_{T},\hat{Q})\in\mathop{\rm argmin}_{(\mathbf{F}_{T},\mathbf{A},q)\in\mathcal{B}_{r}(B)\circ\mathcal{B}_{dd^{*}}(B)\circ\mathcal{Q}}\hat{\mathcal{L}}(\hat{\mathbf{h}},Q_{q,\mathbf{A}},\mathbf{F}_{T})$ (13)
$=$ $\sum_{i=1}^{m}\ell(\mathbf{F}_{T}\hat{\mathbf{h}}(\mathbf{X}_{T,i})+q(\mathbf{A}\mathbf{X}_{T,i}),Y_{T,i})/m+\kappa\hat{\mathcal{V}}_{m}[\hat{\mathbf{h}},Q_{q,\mathbf{A}}]+\chi\|\mathbf{F}_{T}\|_{1}+{\zeta\|\mathbf{A}\|_{F}^{2}},$
where $Q_{q,\mathbf{A}}(\cdot)=q(\mathbf{A}\cdot)$ and $\ell(\cdot,\cdot)$ is the loss for regression or classification, $\mathcal{B}_{r}(B)$ is the ball in $\mathbb{R}^{r}$ with radius $B$, and $\mathcal{Q}$ is a prescribed neural network class, and the empirical distance covariance term $\hat{\mathcal{V}}_{m}[\hat{\mathbf{h}},Q_{q,\mathbf{A}}]$ enforces the independent constraint in (6), and $\kappa,\chi,\zeta$ are the regularization parameters. Also, define the population risk in downstream as $\mathcal{L}(\mathbf{h},Q_{q,\mathbf{A}},\mathbf{F}_{T})=\mathbb{E}_{X_{T},Y_{T}}[\hat{\mathcal{L}}(\mathbf{h}(X_{T}),q(\mathbf{A}X_{T}),\mathbf{F}_{T})]$.
2.2 function class and measure
2.2.1 Distance covariance
We recall the concept of distance covariance [44], which characterizes the dependence of two random variables. Let $\mathfrak{i}$ be the imaginary unit $(-1)^{1/2}$. For any $\mathbf{t}\in\mathbb{R}^{d}$ and $\mathbf{s}\in\mathbb{R}^{m}$, let $\psi_{Z}(\mathbf{t})=\mathbb{E}[\exp^{\mathfrak{i}\mathbf{t}^{T}Z}],\psi_{Y}(\mathbf{s})=\mathbb{E}[\exp^{\mathfrak{i}\mathbf{s}^{T}Y}],$ and $\psi_{Z,Y}(\mathbf{t},\mathbf{s})=\mathbb{E}[\exp^{\mathfrak{i}(\mathbf{t}^{T}Z+\mathbf{s}^{T}Y)}]$ be the characteristic functions of random vectors $Z\in\mathbb{R}^{d},Y\in\mathbb{R}^{m},$ and the pair $(Z,Y)$, respectively. The squared distance covariance $\mathcal{V}[Z,Y]$ is defined as
$\mathcal{V}[Z,Y]=\int_{\mathbb{R}^{d+m}}\frac{\left|\psi_{Z,Y}(\mathbf{t},\mathbf{s})-\psi_{Z}(\mathbf{t})\psi_{Y}(\mathbf{s})\right|^{2}}{c_{d}c_{m}\|\mathbf{t}\|^{d+1}\|\mathbf{s}\|^{m+1}}\mathrm{d}\mathbf{t}\mathrm{d}\mathbf{s},$
where $c_{d}=\frac{\pi^{(d+1)/2}}{\Gamma((d+1)/2)}.$ Given $n$ i.i.d copies $\{Z_{i},Y_{i}\}_{i=1}^{n}$ of $(Z,Y)$, an unbiased estimator of $\mathcal{V}$ is the empirical distance covariance $\widehat{\mathcal{V}}_{n}$, which can be elegantly expressed as a $U$-statistic [45]
$\widehat{\mathcal{V}}_{n}[Z,Y]=\frac{1}{C_{n}^{4}}\sum_{1\leq i_{1}<i_{2}<i_{3}<i_{4}\leq n}k\left(\left(Z_{i_{1}},Y_{i_{1}}\right),\cdots,\left(Z_{i_{4}},Y_{i_{4}}\right)\right),$
where $k$ is the kernel defined by
$k\left(\left(\mathbf{z}_{1},\mathbf{y}_{1}\right),\ldots,\left(\mathbf{z}_{4},\mathbf{y}_{4}\right)\right)$ $=$ $\frac{1}{4}\sum_{\begin{subarray}{c}1\leq i,j\leq 4\\ i\neq j\end{subarray}}\|\mathbf{z}_{i}-\mathbf{z}_{j}\|\|\mathbf{y}_{i}-\mathbf{y}_{j}\|+\frac{1}{24}\sum_{\begin{subarray}{c}1\leq i,j\leq 4\\ i\neq j\end{subarray}}\left\|\mathbf{z}_{i}-\mathbf{z}_{j}\right\|\sum_{\begin{subarray}{c}1\leq i,j\leq 4\\ i\neq j\end{subarray}}\|\mathbf{y}_{i}-\mathbf{y}_{j}\|$
-- $\frac{1}{4}\sum_{i=1}^{4}(\sum_{\begin{subarray}{c}1\leq j\leq 4\\ j\neq i\end{subarray}}\left\|\mathbf{z}_{i}-\mathbf{z}_{j}\right\|\sum_{\begin{subarray}{c}1\leq j\leq 4\\ i\neq j\end{subarray}}\|\mathbf{y}_{i}-\mathbf{y}_{j}\|).$
2.2.2 function class and norm-constrained neural network
Here, we assume that all optimal funcitons are $\beta$-Hölder smooth functions as stated in assumption 1 and 3 below. We aim to develop an approximation theory by utilizing the smoothness of optimal functions and obtain an explicit approximation error bound in terms of the network norm constraints.
Hölder class. Let $\beta=s+r>0,r\in(0,1]$ and $s=\lfloor\beta\rfloor\in\mathbb{N}_{0}$, where $\lfloor\beta\rfloor$ denotes the largest integer strictly smaller than $\beta$ and $\mathbb{N}_{0}$ denotes the set of non-negative integers. For a finite constant $B>0$, the Hölder class of functions $\mathcal{H}^{\beta}\left([0,1]^{d},B\right)$ is defined as
$\mathcal{H}^{\beta}([0,1]^{d},B)=\{f:[0,1]^{d}\rightarrow\mathbb{R},\max_{\|\alpha\|_{1}\leq s}\left\|\partial^{\alpha}f\right\|_{\infty}\leq B,\max_{\|\alpha\|_{1}=s}\sup_{x\neq y}\frac{\left|\partial^{\alpha}f(\mathbf{X})-\partial^{\alpha}f(y)\right|}{\|x-y\|^{r}}\leq B\}$
where $\partial^{\alpha}=\partial^{\alpha_{1}}\cdots\partial^{\alpha_{d}}$ with $\alpha=\left(\alpha_{1},\ldots,\alpha_{d}\right)^{\top}\in\mathbb{N}_{0}^{d}$ and $\|\alpha\|_{1}=\sum_{i=1}^{d}\alpha_{i}$. The Lipschitz class $\mathcal{H}^{1}:=\mathcal{H}([0,1]^{d},K_{2})$ is a special case of Hölder class with $\beta=1$. We adopt the Hölder class as the evaluation class, which encompasses the true functions. This is a common setup in nonparametric estimation [46].
Norm-constrained neural network. We consider a feedforward neural network (FNN) with a ReLU activation function to approximate the functions in the Hölder class. A neural network function $\phi:\mathbb{R}^{N_{0}}\to\mathbb{R}^{N_{L+1}}$ is a function that can be parameterized as,
$\phi(\mathbf{X})=T_{L}(\sigma(T_{L-1}(\cdots\sigma(T_{0}(\mathbf{X}))\cdots))),$ (14)
where the activation function $\sigma(x):=x\lor 0$ is applied component-wisely and $T_{l}(\mathbf{X}):=A_{l}x+b_{l}$ is an affine transformation.
We first introduce a special class of neural network functions $\mathcal{SNN}(W,L)$ which contains functions of the form
$\hat{\phi}(\bm{x})=\hat{A}_{L}\sigma\left(\hat{A}_{L-1}\sigma\left(\cdots\sigma\left(\hat{A}_{0}\hat{\bm{x}}\right)\right)\right),\hat{\bm{x}}:=\left(\begin{array}[]{l}\bm{x}\\ 1\end{array}\right),$ (15)
where $\hat{A}_{\ell}\in\mathbb{R}^{\hat{N}_{\ell+1}\times\hat{N}_{\ell}}$ with $\hat{N_{0}}=d+1$ and $\max\left\{\hat{N}_{1},\ldots,\hat{N}_{L}\right\}=W$. Since these functions can also be written in the form (14) with $\bm{b}_{\ell}=\mathbf{0}$ for all $11\leq\ell\leq L-$ 1, it is obvious that $\mathcal{SNN}(W,L)\subseteq\mathcal{NN}(W,L)$. There is a natural way to introduce norm constraint on the weights: for any $K\geq 0$, we denote by $\mathcal{SN\mathcal{N}}(W,L,K)$ the set of functions in the form (15) that satisfies
$\prod_{\ell=0}^{L}\left\|\hat{A}_{\ell}\right\|\leq K.$
To introduce norm constraint for the class $\mathcal{NN}(W,L)$, we observe that any $\phi\in\mathcal{NN}(W,L)$ parameterized as (14) with $b_{L}=0$ can be written in the form
$\tilde{\phi}(\bm{x})=\tilde{A}_{L}\sigma\left(\tilde{A}_{L-1}\sigma\left(\cdots\sigma\left(\tilde{A}_{0}\tilde{\bm{x}}\right)\right)\right),\quad\tilde{\bm{x}}:=\left(\begin{array}[]{l}\bm{x}\\ 1\end{array}\right)$
with
$\tilde{A}_{L}=\left(A_{L},\mathbf{0}\right),\quad\tilde{A}_{\ell}=\left(\begin{array}[]{cc}A_{\ell}&\bm{b}_{\ell}\\ \mathbf{0}&1\end{array}\right),\ell=0,\ldots,L-1,$
and
$\prod_{\ell=0}^{L}\left\|\tilde{A}_{\ell}\right\|=\left\|A_{L}\right\|\prod_{\ell=0}^{L-1}\max\left\{\left\|\left(A_{\ell},\bm{b}_{\ell}\right)\right\|,1\right\}.$
Hence, we define the norm constrained neural network $\mathcal{NN}(W,L,K)$ as the set of functions $\phi_{\theta}\in\mathcal{NN}(W,L)$ of the form (14) that satisfies the following norm constraint on the weights
$\kappa(\theta):=\left\|A_{L}\right\|\prod_{\ell=0}^{L-1}\max\left\{\left\|\left(A_{\ell},\bm{b}_{\ell}\right)\right\|,1\right\}\leq K.$
We denote by $\mathcal{N}\mathcal{N}_{d_{1},d_{2}}(W,L)$ the set of ReLU FNN functions with width at most $W$, depth at most $L$ and its input (output) dimensions of $d_{1}$($d_{2}$). When the input dimension and output dimension are clear from contexts, we simply denote it by $\mathcal{N}\mathcal{N}(W,L)$. It is desirable to ensure bounded Lipschitz constant of neural networks for ease of analyse. It’s easy to see that any $f\in\mathcal{N}\mathcal{N}(W,L,K)$, $\|f\|_{L}<K$.
We set the network class used in Section 2.1.2-2.1.3 as follows: $\mathcal{H}=\mathcal{N}\mathcal{N}_{d,r}(W_{1},L_{1},K_{1})$, $\mathcal{G}=\mathcal{N}\mathcal{N}_{r,1}(W_{2},L_{2},K_{2})$ and $\mathcal{Q}=\mathcal{N}\mathcal{N}_{d^{*},1}(W_{3},L_{3},K)$.
3 Main Results
3.1 Result for upstream training
We first make the following regularity assumptions on the target function and on the neural network function classes.
Assumption 1.
(Regularity conditions on excess risk).
  • (A1)
    $\mathcal{X}=[0,1]^{d}$, and $|Y|\leq B$ for all $Y\in\mathcal{Y}$;
  • (A2)
    The sparse vector $\|\mathbf{F}^{*}_{s}\|\leq B,\forall s\in\mathcal{S}$, and each component of $\mathbf{h}^{*}$ is contained in $\mathcal{H}^{\beta}([0,1]^{d},B)$;
  • (A3)
    $\sup_{\mathbf{x}\in\mathcal{X}}\{\|g(\mathbf{h}(\mathbf{x}))\|\lor\|\mathbf{h}(\mathbf{x})\|\}\leq B$ for all $g\in\mathcal{G},\mathbf{h}\in\mathcal{H}$.
Remark 2.
The regularity conditions in Assumption 1 are mild. Firstly, we impose the boundedness assumptions on $\mathcal{X}$ and $\mathcal{Y}$ in (A1) for the sake of simplicity in presentation. This choice is facilitated by the ability to employ truncation techniques, given additional exponential decay tail assumptions on the distribution of the covariate and the noise, as elaborated in detail in [46]. Secondly, the assumption of a bounded target (A2) is standard in both parametric and nonparametric estimation. We required the network’s output to be bounded in (A3), which is achievable through a truncation implemented by a ReLU layer.
Theorem 1 (Upstream excess risk bound).
Suppose the Assumption 1 holds. Set $\mathcal{H}=\mathcal{N}\mathcal{N}_{d,r}(W_{1},L_{1},K_{1})$ and network $\mathcal{G}=\mathcal{N}\mathcal{N}_{r,1}(W_{2},L_{2},K_{2})$ according the following Table:
NN class $\mathcal{H}$ $\mathcal{G}$
width(W) $W_{1}\asymp n^{\frac{2d+\beta}{4(d+1+\beta)}}\mathbb{I}_{\beta\leq 2}+n^{\frac{2d+\beta}{2(2d+3\beta)}}\mathbb{I}_{\beta>2}$ $W_{2}\gtrsim n^{\frac{2r+\beta}{4(d+1+\beta)}}\mathbb{I}_{\beta\leq 2}+n^{\frac{2r+\beta}{2(2d+3\beta)}}\mathbb{I}_{\beta>2}$
depth(L) $L_{1}\geq 2\left\lceil\log_{2}(r+\lfloor\beta\rfloor)\right\rceil+2$ $L_{2}\geq 2\left\lceil\log_{2}(r+\lfloor\beta\rfloor)\right\rceil+2$
norm constraint(K) $K_{1}\asymp n^{\frac{d+1}{2(d+\beta+1)}}\mathbb{I}_{\beta\leq 2}+n^{\frac{d+1}{2d+3\beta}}\mathbb{I}_{\beta>2}$ $K_{2}\asymp n^{\frac{r+1}{2(d+1+\beta)}}\mathbb{I}_{\beta\leq 2}+n^{\frac{r+1}{2d+3\beta}}\mathbb{I}_{\beta>2}$
Then the ERM solution $(\hat{\mathbf{F}},\hat{\mathbf{h}})$ from (12) satisfy:
$\mathbb{E}[R(\hat{\mathbf{F}},\hat{\mathbf{h}})-R(\mathbf{F}^{*},\mathbf{h}^{*})]\precsim\tilde{O}\left(n^{-\frac{\beta}{2(d+1+\beta)}}\mathbb{I}_{\beta\leq 2}+n^{-\frac{\beta}{2d+3\beta}}\mathbb{I}_{\beta>2}\right).$ (16)
Remark 3.
Theorem 1 illustrates that, with an appropriate configuration of neural networks in terms of width, depth, and norm constraints, it is possible to design ERM solutions for upstream representation learning that exhibit a rate for nonparametric estimation.
A key feature of Theorem 1 is that, by appropriately controlling the norm constraint on the network, the rate can be achieved with only the restriction on the lower bound of the depth and width for $\mathcal{G}$. Hence, we can attain the consistency results within an over-parameterized network class, representing a novel finding in the realm of deep transfer learning, to the best of our current understanding.
Tripuraneni et al. [21, Theorem 4] presented a generalization bound for multitask transfer learning. However, their analysis is based on a well-specified neural network model, where the target function is also a network, without accounting for the approximation error as we have done.
Since the true vector $\mathbf{F}^{*}_{s},s\in\mathcal{S}$ is sparse, we also want to derive the consistency for the estimation error. Hence, we introduce the following assumptions.
Assumption 2.
(Regularity conditions on estimation).
  • (A4)
    Denote $\hat{\mathbf{H}}_{i}=\frac{1}{\sqrt{n}}(\hat{\mathbf{h}}(X_{i,1}),\ldots,\hat{\mathbf{h}}(X_{i,n_{i}}))$. We assume that $\|\hat{\mathbf{H}}_{i}\|_{2}>\sqrt{B_{1}}$ for all the $i\in[p]$, where $B_{1}>0$ is a constant;
  • (A5)
    The ERM solution $\hat{\mathbf{h}}$ from (12) satisfied $\mathbb{E}[\|\mathbf{h}^{*}-\hat{\mathbf{h}}\|_{L_{1}(\mathbb{P}_{\mathbf{X}})}]\asymp\mathbb{E}[\mathcal{W}_{1}(\hat{\mathbf{h}}(\mathbf{X}),\mathbf{h}^{*}(\mathbf{X}))]$.
Remark 4.
Assumption (A4) is mild due to the condition $r\ll n$, ensuring that each $\hat{\mathbf{H}}_{i}^{T}$ easily becomes full column rank. Assumption (A5) is a technical condition which requires that there exist a constant $B_{2}$ such that $\mathbb{E}[\|\mathbf{h}^{*}-\hat{\mathbf{h}}\|_{L_{1}(\mathbb{P}_{\mathbf{X}})}]\leq B_{2}\mathbb{E}[\mathcal{W}_{1}(\hat{\mathbf{h}}(\mathbf{X}),\mathbf{h}^{*}(\mathbf{X}))]$, given another direction of the inequality holds trivially.
Theorem 2 (Estimator error of $\hat{\mathbf{F}}$).
Under Theorem 1 and Assumption 2, for any $\delta>0$, there exits constant $\gamma>0$ and a constant $C(\gamma,B,B_{1},B_{2},r,p,\delta)$ such that if we set $\mu=\frac{\gamma}{2Brp}(n^{-\frac{\beta}{2(d+1+\beta)}}\mathbb{I}_{\beta\leq 2}+n^{-\frac{\beta}{2d+3\beta}}\log n\mathbb{I}_{\beta>2})$, then with probability at least $1-2\delta$ over the draw of dataset $\mathcal{D}$,
$\|\hat{\mathbf{F}}-\mathbf{F}^{*}\|_{2}\leq C(\gamma,B,B_{1},B_{2},r,p,\delta)(n^{-\frac{\beta}{2(d+1+\beta)}}\mathbb{I}_{\beta\leq 2}+n^{-\frac{\beta}{2d+3\beta}}\log n\mathbb{I}_{\beta>2}).$
3.2 Results for downstream prediction
First, we give the assumptions on the smoothness of target function, and transferable condition.
Assumption 3.
(Regularity conditions for down-stream task). (A6) Each entries of $q^{*}$ is contained in $\mathcal{H}^{\beta}([-1,1]^{d},B)$. $\|\mathbf{F}_{T}^{*}\|_{\infty}\leq B$ and $\|\mathbf{A}^{*}\|_{F}\leq B$; (A7) $\ell(\cdot,y)$ is Lipschitz with respect to $y\in\mathcal{Y}$ and the Lipschitz constant is bounded by $L_{0}$, and $|\ell(\cdot,\cdot)|\leq B_{3}$; (A8) $\mathcal{W}_{1}(\mathbf{X},\mathbf{X}_{T})\leq\omega$ with $\omega\asymp\left(n^{-\frac{1}{2}}\mathbb{I}_{\beta\leq 2}+n^{-\frac{\beta+d+1}{2d+3\beta}}\mathbb{I}_{\beta>2}\right)$; (A9) Consider $\mathcal{H}_{\eta}=\{h\in\mathcal{H}|\mathbb{E}[\mathcal{W}_{1}(\mathbf{h}(X_{T}),\mathbf{h}^{*}(X_{T}))]\leq\eta,\eta\precsim\tilde{O}(n^{-\frac{\beta}{2(d+1+\beta)}}\mathbb{I}_{\beta\leq 2}+n^{-\frac{\beta}{2d+3\beta}}\mathbb{I}_{\beta>2})\},$ suppose $\sup_{\mathbf{h}\in\mathcal{H}_{\delta}}\mathbb{E}[\mathbb{E}_{Y_{T}}[\mathcal{W}_{1}(\mathbf{F}_{T}^{*}\mathbf{h}(X_{T})|Y_{T},\mathbf{F}_{T}^{*}\mathbf{h}^{*}(X_{T})|Y_{T})]]\leq\nu\mathbb{E}[\mathcal{W}_{1}(\mathbf{h}(X_{T}),\mathbf{h}^{*}(X_{T}))],$ where $\nu>0$ is a constant.
Remark 5.
Conditions (A6)-(A7) are standard requirements in the context of nonparametric supervised learning. Condition (A8) characterizes the similarity between the upstream and downstream tasks, a prerequisite for transfer feasibility. Let $\mathbb{P}_{\mathbf{X}}$ and $\mathbb{P}_{T}$ denote the marginal distributions of $\mathbf{X}$ and $X_{T}$ in the upstream and downstream tasks, respectively. In previous work, specifically in [21], the condition $\mathbb{P}_{\mathbf{X}}=\mathbb{P}_{T}$ or $\mathcal{W}_{1}(\mathbf{X},X_{T})=0$ was deemed necessary. In contrast, our assumption relaxes this requirement, allowing $\mathcal{W}_{1}(\mathbf{X},X_{T})$ to be small, indicating that a minor shift in input data is acceptable. Condition (A9) constitutes a technical requirement on local conditional distribution matching. Specifically, it implies that when the distribution of $\mathbf{h}(X_{T})$ closely approximates that of $\mathbf{h}^{*}(X_{T})$, the average conditional Wasserstein distance between the representations $\mathbf{h}(X_{T})$ and $\mathbf{h}^{*}(X_{T})$ is also small. In a recent study on domain adaptation, a similar, albeit more stringent, assumption known as conditional invariant component is employed, as outlined in [47].
Theorem 3 (Excess risk bound for downstream task).
Under the Assumptions 1, 3, we plug the learned representation $\hat{\mathbf{h}}$ described in Theorem 1 into (13), and set the network class $\mathcal{Q}=\mathcal{NN}_{d^{*},1}(W_{3},L_{3},K)$ satisfying $W_{3}\gtrsim m^{\frac{2d^{*}+\beta}{4(d^{*}+1+\beta)}},L_{3}\geq 2\left\lceil\log_{2}(d^{*}+s)\right\rceil+2,K\asymp m^{\frac{d^{*}+1}{2(d^{*}+1+\beta)}}.$ Then ERM solution $\hat{\mathbf{F}}_{T}$ and $\hat{Q}$ enjoy the following
$\mathbb{E}[\mathcal{L}(\hat{h},\hat{Q},\hat{\mathbf{F}}_{T})-\mathcal{L}(\mathbf{h}^{*},Q^{*},\mathbf{F}^{*}_{T})]\leq\tilde{O}\left(n^{-\frac{\beta}{2(d+1+\beta)}}\mathbb{I}_{\beta\leq 2}+n^{-\frac{\beta}{2d+3\beta}}\mathbb{I}_{\beta>2}\right)+\mathcal{O}\left(m^{-\frac{\beta}{2(d^{*}+1+2\beta)}}\right).$
Remark 6.
The above Theorem 3 demonstrates that end-to-end prediction accuracy for the downstream task is $\mathcal{O}(m^{-\frac{\beta}{2(d^{*}+1+2\beta)}})$ if the error in upstream training is negligible due to the huge sample size $n$.
Remark 7.
In the challenging scenario of “no transfer” where $\mathbf{F}_{T}^{*}=0$ and $d^{*}=d$, the convergence rate goes back to $\mathcal{O}(m^{-\frac{\beta}{2(d+1+2\beta)}})$. This rate aligns with the convergence rate achieved using ERM for learning the $d$-dimensional function $Q^{*}$ in classification [48] and is slower than the minimax rate observed in regression [49, 50, 51, 52, 53, 54, 55, 56]. Again, same with Remark 3 on Theorem 1, the convergence rate established in Theorem 3 is applicable to deep learning models characterized by over-parametrization. In the case of “partial transfer”, where $\mathbf{F}_{T}^{*}\neq 0$ and $Q^{*}(\cdot)=q^{*}(\mathbf{A}^{*}\cdot)$ with $d^{*}<d$, the convergence rate demonstrates improvement from $\mathcal{O}(m^{-\frac{\beta}{2(d+1+2\beta)}})$ to $\mathcal{O}(m^{-\frac{\beta}{2(d^{*}+1+2\beta)}})$, benefiting from the knowledge embedded in the upstream representation. In the case of “complete transfer” where $Q^{*}=0$, the convergence rate experiences further enhancement from $\mathcal{O}(m^{-\frac{\beta}{2(d+1+2\beta)}})$ to $\mathcal{O}(\frac{1}{\sqrt{m}})$. This enhancement is elaborated in detail in the following Theorem 4. Before this, we first give some assumptions.
Assumption 4.
(Regularity conditions on estimation error for downstream task).
  • (A10)
    Denote $\hat{\mathbf{H}}_{T}=\frac{1}{\sqrt{m}}(\hat{\mathbf{h}}(\mathbf{X}_{T,1}),\ldots,\hat{\mathbf{h}}(\mathbf{X}_{T,m}))$. We assume that $\|\hat{\mathbf{H}}_{T}\|_{2}>\sqrt{B_{1}}$;
  • (A11)
    $\mathbb{E}[\|\mathbf{h}^{*}-\hat{\mathbf{h}}\|_{L_{1}(\mathbb{P}_{X_{T}})}]\asymp\mathbb{E}[\mathcal{W}_{1}(\hat{\mathbf{h}}(X_{T}),\mathbf{h}^{*}(X_{T}))]$.
Theorem 4.
(Estimation error for “complete transfer”) Under the Assumptions 1, 3 and 4, for any $\delta>0$, there exits constant $\gamma_{1}>0$, constant $C_{T,1}(\gamma_{1},B,B_{2},r,\delta)$ and $C_{T,2}(\gamma_{1},B,B_{2},r,\delta)$ such that if we set $\chi=\frac{1}{\sqrt{m}}$ in the least squares regression for complete transfer, then with probability at least $1-2\delta$ over the draw of dataset $\mathcal{T}$
$\mathbb{E}_{\mathcal{D}}[\|\mathbf{F}^{*}_{T}-\hat{\mathbf{F}}_{T}\|_{2}]$
$~{}~{}~{}~{}\leq C_{T,1}(B,r,\delta)m^{-\frac{1}{2}}+C_{T,2}(\gamma_{1},B,B_{2},r,\delta)(n^{-\frac{\beta}{2(d+1+\beta)}}\mathbb{I}_{\beta\leq 2}+n^{-\frac{\beta}{2d+3\beta}}\log n\mathbb{I}_{\beta>2}).$
It shows the sparsity of $\hat{\mathbf{F}}_{T}$ when assumption 3 and 4 holds and m is considerable large.
Remark 8.
Notably, in the binary classification problem, we can change the loss into logistic loss and utilize its local strongly convexity to get the similar result.
4 Experiment
We substantiate the effectiveness of our method through experimental validation using four image classification datasets and one regression dataset. More details on datasets and hyperparameter setting are shown in Appendix E.
Datasets. The four classification datasets are VLCS [57], PACS [58], TerraIncognita [59] and OfficeHome [60]. All these datasets comprise image data with four distinct domains. The regression dataset is the tool wear dataset from the 2010 PHM Challenge [61], which comprises labeled data with three domains, namely C1, C4, and C6. The original dataset consists of sensing signals sampled at 50 kHz, including cutting force signals, vibration signals in the x, y, and z directions, and acoustic emission signals during processing. We use cutting force signals in three directions as input and the average wear values of the three blades for each tool as output. Since the results for the tool wear were not significant across all domains, we have presented them in Appendix E.
Model selection. In the classification task, we partition data from different domains into training and validation sets in an 8:2 ratio. The transfer process involves the following steps: withholding data from one domain, merging the training sets of the remaining domains to construct the upstream training set and the validation sets to construct the validation set accordingly. For each domain, we vary the hyperparameter $\lambda$, $\tau$ and $\mu$, selecting the model with the highest accuracy on the validation set we construct from the upstream training set. During the downstream task, training is conducted utilizing the training set from the reserved domain, with adjusting the hyperparameter $\kappa,\chi,\zeta$ and $d^{*}$. The model with the highest training accuracy is chosen, and its accuracy on the withheld domain’s validation set is considered as the final accuracy. In the regression task, the distinction lies in partitioning data from various domains into training and validation sets at a 9:1 ratio. Meanwhile, Mean Absolute Error (MAE), Root Mean Square Error (RMSE), and Coefficient of determination (R2) [62] are used to evaluate performance of methods.
Table 2: The accuracy of different training methods on typical datasets and their domains
Dataset/Algorithm Accuracy on domain
officehome A C P R Avg
ERM-UD 71.4 $\pm$ 1.0 73.3 $\pm$ 0.1 90.3 $\pm$ 0.1 83.8 $\pm$ 0.6 79.7 $\pm$ 0.1
TIR 71.1 $\pm$ 0.6 72.9 $\pm$ 0.6 89.8 $\pm$ 0.7 82.7 $\pm$ 0.3 79.1 $\pm$ 0.1
ERM-D 51.2 $\pm$ 1.9 60.4 $\pm$ 1.2 83.1 $\pm$ 0.3 78.8 $\pm$ 0.5 68.4 $\pm$ 0.2
WI 70.6 $\pm$ 0.7 73.3 $\pm$ 0.1 89.7 $\pm$ 0.4 83.2 $\pm$ 0.5 79.2 $\pm$ 0.3
Ours 77.5 $\pm$ 1.1 78.4 $\pm$ 0.3 93.3 $\pm$ 0.1 86.7 $\pm$ 0.5 84.0 $\pm$ 0.3
PACS A C P S Avg
ERM-UD 92.3 $\pm$ 0.7 93.7 $\pm$ 0.6 99.0 $\pm$ 0.3 88.8 $\pm$ 0.5 93.5 $\pm$ 0.2
TIR 94.2 $\pm$ 0.2 94.4 $\pm$ 0.1 99.2 $\pm$ 0.2 89.6 $\pm$ 0.4 94.3 $\pm$ 0.1
ERM-D 89.5 $\pm$ 0.2 88.4 $\pm$ 0.3 99.1 $\pm$ 0.3 90.8 $\pm$ 0.3 92.0 $\pm$ 0.2
WI 94.2 $\pm$ 0.2 93.6 $\pm$ 0.2 98.9 $\pm$ 0.5 89.3 $\pm$ 0.6 94.0 $\pm$ 0.1
Ours 95.0 $\pm$ 0.4 96.0 $\pm$ 0.4 99.4 $\pm$ 0.3 92.6 $\pm$ 0.7 95.7 $\pm$ 0.4
VLCS C L S V Avg
ERM-UD 99.4 $\pm$ 0.2 72.2 $\pm$ 2.0 78.2 $\pm$ 1.0 84.0 $\pm$ 0.4 83.5 $\pm$ 0.7
TIR 99.8 $\pm$ 0.2 74.8 $\pm$ 1.0 80.6 $\pm$ 1.0 83.6 $\pm$ 0.9 84.7 $\pm$ 0.6
ERM-D 100.0 $\pm$ 0.0 73.0 $\pm$ 1.0 79.8 $\pm$ 0.4 86.9 $\pm$ 0.2 85.0 $\pm$ 0.4
WI 99.5 $\pm$ 0.2 74.8 $\pm$ 0.4 80.2 $\pm$ 0.5 84.6 $\pm$ 0.6 84.8 $\pm$ 0.1
Ours 100.0 $\pm$ 0.0 73.9 $\pm$ 0.0 80.0 $\pm$ 0.1 87.3 $\pm$ 0.3 85.3 $\pm$ 0.1
TerraInc 38 43 46 100 Avg
ERM-UD 85.2 $\pm$ 0.4 82.1 $\pm$ 0.6 74.2 $\pm$ 0.5 89.3 $\pm$ 0.6 82.7 $\pm$ 0.3
TIR 83.1 $\pm$ 0.8 76.8 $\pm$ 0.9 72.9 $\pm$ 0.8 86.2 $\pm$ 0.7 79.7 $\pm$ 0.6
ERM-D 85.1 $\pm$ 0.4 69.9 $\pm$ 0.3 70.6 $\pm$ 0.5 83.9 $\pm$ 0.3 77.5 $\pm$ 0.1
WI 82.2 $\pm$ 0.6 76.9 $\pm$ 0.6 71.7 $\pm$ 0.8 85.8 $\pm$ 0.7 79.2 $\pm$ 0.3
Ours 87.4 $\pm$ 0.4 83.5 $\pm$ 0.6 76.3 $\pm$ 0.3 89.6 $\pm$ 0.4 84.2 $\pm$ 0.2
[…]
Comparative methods and their results. We present the test results in Table 2 for the following training methods. ERM-UD (ERM on both upstream and downstream): Training a network without considering task diversity, transferring the feature extractor to the downstream task, and training a linear classifier while freezing the feature extractor. Transfer Invariant Representation (TIR): Adopting and freezing the feature extractor using our upstream training method, followed by training a linear classifier at the downstream task without the $Q$ network. ERM-D (ERM only at downstream): Solely utilizing the pretrained EfficientNet to train the downstream task. Without Independence (WI): Adopting the feature extractor through our upstream training method and incorporating the pretrained EfficientNet while training at the downstream task without imposing any independence constraint between the output of EfficientNet and the transfer featurizer.
Results interpretation. Contrasting the results between our method and WI demonstrates the efficacy of the independence penalty term in improving accuracy for downstream tasks, which can be attributed to that the independence constraints (6) and (7) reduce the learning difficulty of the downstream task. Moreover, when comparing the outcomes of our approach with those of ERM-D, it becomes evident that the enhanced accuracy cannot be solely attributed to the information learned in $Q$. Furthermore, the comparison between TIR and WI reveals that the auxiliary network for extracting different information can be detrimental to accuracy when independence requirements are not imposed. These numerical results align consistently with our observations in Theorem 3 and Remark 6. Specifically, the model under “partial transfer” conditions is demonstrated to be more easily learnable, primarily attributed to not only the advantageous representation obtained from the upstream domains but also the imposition of the independence constraint, which significantly reduces the estimation complexity in the downstream task.
Comparison with other works. While our research setting diverges from that of Domain Adaptation, Domain Generalization, and Multi-task Learning as discussed in [66]. And it also differs from single-source domain transfer learning, as discussed in [4]. We conducted comparisons among several Domain Adaptation methods, as outlined in Table 3, utilizing the code from DomainBed [67]. Despite these methods do not impose any constraints on downstream training, the test results demonstrate the superiority of our approach to some extent. Further implementation details regarding the comparison are provided in Appendix E.
Table 3: The accuracy of different training methods on typical datasets
Algorithm TerraInc PACS VLCS Office-home Avg
ERM[68] 87.1 $\pm$ 0.1 94.1 $\pm$ 0.3 82.0 $\pm$ 0.6 79.9 $\pm$ 0.3 85.9 $\pm$ 0.3
DRO[69] 87.5 $\pm$ 0.5 94.4 $\pm$ 0.0 82.2 $\pm$ 0.5 80.9 $\pm$ 0.5 85.7 $\pm$ 0.2
IRM[70] 82.2 $\pm$ 0.9 93.6 $\pm$ 0.1 82.0 $\pm$ 0.4 76.2 $\pm$ 0.4 83.5 $\pm$ 0.3
DANN[71] 87.1 $\pm$ 0.2 93.7 $\pm$ 0.4 82.1 $\pm$ 0.9 79.4 $\pm$ 0.1 85.6 $\pm$ 0.3
SagNet[72] 87.3 $\pm$ 0.3 94.0 $\pm$ 0.5 81.7 $\pm$ 0.3 79.5 $\pm$ 0.4 85.7 $\pm$ 0.1
Ours 84.2 $\pm$ 0.2 95.7 $\pm$ 0.4 85.3 $\pm$ 0.1 84.0 $\pm$ 0.3 87.3 $\pm$ 0.1
5 Conclusion
In this paper, we propose a statistical framework for transfer learning from a multi-task upstream scenario to a single downstream task. Our framework accommodates both shared and specific features, explicitly identifying the upstream features contributing to downstream tasks. We introduce a novel training algorithm tailored for transferring a suitably invariant shared representation. Convergence analysis reveals that the derived rate is adaptive to the difficulty of transferability. Empirical experiments on real classification and regression data validate the efficacy of our model and theoretical findings.
6 Acknowledgements
Appendices
Appendix A proof of error decomposition and approximation decomposition
$\mathcal{H}$ to be a function class of features mapping $\mathbb{R}^{d}\to\mathbb{R}^{r}$, $\mathcal{G}$ to be a function class of lipschitz-constrained mapping $\mathbb{R}^{r}\to\mathbb{R}$ and $\mathcal{Q}:\mathbb{R}^{d^{*}}\to\mathbb{R}$ to be a function class during fine-tuning. Also, for any $i\in[p]$, $\mathbb{P}(S=i)>0$, which is a general case and means $n_{i}$ is the same order of n. Without loss of generality, we set the upper bound of random noise as 1: $|\varepsilon_{i}|\leq 1,i\in[2]$.
The proof and definitions are organized as follows: In the Section B, we present the proof of Theorem 1 and Theorem 2 in Section 3.1. Section C is dedicated to proving the Theorem 3 and Theorem 4 in Section 3.2. Section D compiles a list of the theorems cited, and Section E provides additional experimental details.
Appendix B Proofs of Section 3.1
For all the following lemmas, we omit the statement about assumptions for simplicity. The lemma proof in Section B.1.1 and B.1.2 is under the assumption we made in Theorem 1. The lemma proof in Section B.2 is under the assumption we made in Theorem 2, the lemma proof in downstream is under the assumption we made in Theorem 3.
B.1 Proofs of Theorem 1
B.1.1 Proof of error decomposition and approximation decomposition
Without loss of generality, we set the $\lambda=1$ and $\tau=1$. We first presents a basic inequality for the excess risk in terms of stochastic and approximation errors.
Lemma 1.
$\mathbb{E}[R(\hat{\mathbf{F}},\hat{\mathbf{h}})-R(\mathbf{F}^{*},\mathbf{h}^{*})]\leq$ $2\sup_{(\mathbf{F},\mathbf{h})\in\mathcal{B}_{pr}(B)\circ\mathcal{H}}\mathbb{E}[|\hat{R}(\mathbf{F},\mathbf{h})-R(\mathbf{F},\mathbf{h})|]$
$+\inf_{(\mathbf{F},\mathbf{h})\in\mathcal{B}_{pr}(B)\circ\mathcal{H}}|R(\mathbf{F},\mathbf{h})-R(\mathbf{F}^{*},\mathbf{h}^{*})|.$
Proof.
For any $\mathbf{F}\in\mathcal{B}_{pr}(B)$ and any $\mathbf{h}\in\mathcal{H}$, we have:
$\mathbb{E}[R(\hat{\mathbf{F}},\hat{\mathbf{h}})-R(\mathbf{F}^{*},\mathbf{h}^{*})]=$ $\mathbb{E}[R(\hat{\mathbf{F}},\hat{\mathbf{h}})-\hat{R}(\hat{\mathbf{F}},\hat{\mathbf{h}})+\hat{R}(\hat{\mathbf{F}},\hat{\mathbf{h}})-\hat{R}(\mathbf{F},\mathbf{h})+\hat{R}(\mathbf{F},\mathbf{h})-R(\mathbf{F},\mathbf{h})$
$+R(\mathbf{F},\mathbf{h})-R(\mathbf{F}^{*},\mathbf{h}^{*})]$
$\leq$ $\mathbb{E}[R(\hat{\mathbf{F}},\hat{\mathbf{h}})-\hat{R}(\hat{\mathbf{F}},\hat{\mathbf{h}})+\hat{R}(\mathbf{F},\mathbf{h})-R(\mathbf{F},\mathbf{h})+R(\mathbf{F},\mathbf{h})-R(\mathbf{F}^{*},\mathbf{h}^{*})]$
$\leq$ $2\sup_{(\mathbf{F},\mathbf{h})\in\mathcal{B}_{pr}(B)\circ\mathcal{H}}\mathbb{E}[|\hat{R}(\mathbf{F},\mathbf{h})-R(\mathbf{F},\mathbf{h})|]+\mathbb{E}[R(\mathbf{F},\mathbf{h})-R(\mathbf{F}^{*},\mathbf{h}^{*})].$
The term $\hat{R}(\hat{\mathbf{F}},\hat{\mathbf{h}})-\hat{R}(\mathbf{F},\mathbf{h})<0$ by definition. Due to the randomness of $\mathbf{F}$ and $\mathbf{h}$, we can take the infimum of both the LHS and the RHS of the inequality. Thus,
$\mathbb{E}[R(\hat{\mathbf{F}},\hat{\mathbf{h}})-R(\mathbf{F}^{*},\mathbf{h}^{*})]\leq$ $2\sup_{(\mathbf{F},\mathbf{h})\in\mathcal{B}_{pr}(B)\circ\mathcal{H}}\mathbb{E}[|\hat{R}(\mathbf{F},\mathbf{h})-R(\mathbf{F},\mathbf{h})|]$
$+\inf_{(\mathbf{F},\mathbf{h})\in\mathcal{B}_{pr}(B)\circ\mathcal{H}}|R(\mathbf{F},\mathbf{h})-R(\mathbf{F}^{*},\mathbf{h}^{*})|\negmedspace.$
Lemma 2.
$\inf_{(\mathbf{F},\mathbf{h})\in\mathcal{B}_{pr}(B)\circ\mathcal{H}}|R(\mathbf{F},\mathbf{h})-R(\mathbf{F}^{*},\mathbf{h}^{*})|\leq|R(\mathbf{F}^{*},\mathbf{h})-R(\mathbf{F}^{*},\mathbf{h}^{*})|\leq(4B^{2}+32p+1)\|\mathbf{h}-\mathbf{h}^{*}\|_{\infty}.$ (17)
Proof.
$|R(\mathbf{F}^{*},\mathbf{h})-R(\mathbf{F}^{*},\mathbf{h}^{*})|$ $\leq|\mathrm{E}[(Y_{S}-\mathbf{F}_{S}^{*}\mathbf{h}(\mathbf{X}_{S}))^{2}]+\mathcal{V}[\mathrm{D},\mathbf{h}(X)]-\mathrm{E}[(Y_{S}-\mathbf{F}_{S}^{*}\mathbf{h}^{*}(\mathbf{X}_{S}))^{2}]$
$\quad-\mathcal{V}[\mathrm{D},\mathbf{h}^{*}(\mathbf{X})]|+\mathcal{W}_{1}(\mathbf{h}(X),\mathcal{U}_{r})-\mathcal{W}_{1}(\mathbf{h}^{*}(\mathbf{X}),\mathcal{U}_{r})$
$\leq 4B\mathrm{E}[|\mathbf{F}_{S}^{*}(\mathbf{h}(X))-\mathbf{F}_{S}^{*}(\mathbf{h}^{*}(\mathbf{X}))|]+|\mathcal{V}[\mathrm{D},\mathbf{h}(X)]-\mathcal{V}[\mathrm{D},\mathbf{h}^{*}(\mathbf{X})]|$
$\quad+\mathcal{W}_{1}(\mathbf{h}(X),\mathbf{h}^{*}(\mathbf{X}))$
$\leq 4B^{2}\|\mathbf{h}-\mathbf{h}^{*}\|_{\infty}+|\mathcal{V}[\mathrm{D},\mathbf{h}(X)]-\mathcal{V}[\mathrm{D},\mathbf{h}^{*}(\mathbf{X})]|+\mathcal{W}_{1}(\mathbf{h}(X),\mathbf{h}^{*}(\mathbf{X}))$
$\leq(4B^{2}+32p+1)\|\mathbf{h}-\mathbf{h}^{*}\|_{\infty}.$
Here we utilize the metric property of Wasserstein distance and propery of covariance distance in Székely et al. [44, p16,Remark 3]. ∎
B.1.2 Statistical errors
We can decompose the analysis of statistical errors into three components by the loss we defined.
$\mathbb{E}[\sup_{(\mathbf{F},\mathbf{h})\in\mathcal{B}_{pr}(B)\circ\mathcal{H}}|\hat{R}(\mathbf{F},\mathbf{h})-R(\mathbf{F},\mathbf{h})|]$ (18)
$\leq$ $\mathbb{E}[\underbrace{\sup_{(\mathbf{F},\mathbf{h})\in\mathcal{B}_{pr}(B)\circ\mathcal{H}}|\mathbb{E}[(Y_{S}-\mathbf{F}_{S}\mathbf{h}(\mathbf{X}_{S}))^{2}]-\frac{1}{n}\sum_{i=1}^{n}(Y_{S_{i},i}-\mathbf{F}_{S_{i}}(\mathbf{h}(x_{S_{i},i})))^{2}|}_{\mathcal{L}_{1}}$
$+\underbrace{\mathbb{E}[\sup_{\mathbf{h}\in\mathcal{H}}|V[\mathbf{h}(X),D]-\widehat{\mathcal{V}}[\mathbf{h}(X),D]|]}_{\mathcal{L}_{2}}+\underbrace{\mathbb{E}[\sup_{\mathbf{h}\in\mathcal{H}}|\hat{\mathcal{W}_{1}}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})-\mathcal{W}_{1}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})|]}_{\mathcal{L}_{3}}.$
We decompose this part of the error into three components and analyze them separately. The processing for bounding the statistical error of $\mathcal{L}_{1}$ and $\mathcal{L}_{3}$ is almost standard [73, 74]. In the context of analysis restricted to $\mathcal{L}_{2}$, we leveraged the methodological techniques concerning the estimation error of the U-statistic as presented in [75].
In the following we will bounding the three term in (18) one by one.
Lemma 3.
$\mathcal{L}_{1}\leq 8B\mathbb{E}[\sup_{(\mathbf{F},\mathbf{h})\in\mathcal{B}_{pr}(B)\circ\mathcal{H}}|\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\mathbf{F}_{S_{i}}\mathbf{h}(\mathbf{X}_{S_{i},i})|]+\frac{2B^{2}}{\sqrt{n}}.$ (19)
Proof.
Consider ${X^{\prime}_{S^{\prime}_{i},i},Y^{\prime}_{S^{\prime}_{i},i},S^{\prime}_{i}}_{i=1}^{n}$ be a ghost sample of i.i.d random variables distributed as $\mathbb{P}(\mathbf{x},y,s)$ and independent of dataset $\mathcal{D}$. $\sigma_{i}$ is i.i.d random variables distributed as Rademacher distribution $Rad(\frac{1}{2})$. And $\sigma_{i}$ is independent of all other randomness.
$\mathcal{L}_{1}=$ $\mathbb{E}[\sup_{(\mathbf{F},\mathbf{h})\in\mathcal{B}_{pr}(B)\circ\mathcal{H}}|\mathbb{E}[(Y_{S}-\mathbf{F}_{S}\mathbf{h}(\mathbf{X}_{S}))^{2}]-\frac{1}{n}\sum_{i=1}^{n}(Y_{S_{i},i}-\mathbf{F}_{S_{i}}(\mathbf{h}(\mathbf{x}_{S_{i},i})))^{2}|]$
$\leq$ $\mathbb{E}[\sup_{(\mathbf{F},\mathbf{h})\in\mathcal{B}_{pr}(B)\circ\mathcal{H}}|\frac{1}{n}\sum_{i=1}^{n}(Y^{\prime}_{S^{\prime}_{i},i}-\mathbf{F}_{S^{\prime}_{i}}\mathbf{h}(\mathbf{X}^{\prime}_{S^{\prime}_{i},i})))^{2}-\frac{1}{n}\sum_{i=1}^{n}(Y_{S_{i},i}-\mathbf{F}_{S_{i}}\mathbf{h}(\mathbf{X}_{S_{i},i}))^{2}|]$
$\leq$ $\mathbb{E}[\sup_{(\mathbf{F},\mathbf{h})\in\mathcal{B}_{pr}(B)\circ\mathcal{H}}|\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}(Y^{\prime}_{S^{\prime}_{i},i}-\mathbf{F}_{S^{\prime}_{i}}\mathbf{h}(\mathbf{X}^{\prime}_{S^{\prime}_{i},i})))^{2}-\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}(Y_{S_{i},i}-\mathbf{F}_{S_{i}}\mathbf{h}(\mathbf{X}_{S_{i},i}))^{2}|]$
$\leq$ $2\mathbb{E}[\sup_{(\mathbf{F},\mathbf{h})\in\mathcal{B}_{pr}(B)\circ\mathcal{H}}|\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}(Y_{S_{i},i}-\mathbf{F}_{S_{i}}\mathbf{h}(\mathbf{X}_{S_{i},i}))^{2}|]$
$\leq$ $8B\mathbb{E}[\sup_{(\mathbf{F},\mathbf{h})\in\mathcal{B}_{pr}(B)\circ\mathcal{H}}|\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\mathbf{F}_{S_{i}}\mathbf{h}(\mathbf{X}_{S_{i},i})|]+\frac{2B^{2}}{\sqrt{n}}.$
The last inequality is due to the Wainwright [76, Exercise 4.7c] and the contraction principle Ledoux and Talagrand [77, Theorem 4.12]. ∎
As for the upper bound of $\mathcal{L}_{2}$, We need an alternative pseudo-metric to simplify the analysis. $\forall\mathbf{h}\in\mathcal{H}$, define a random empirical measure (depends on $\tilde{D}_{i}=(\mathbf{X}_{i},S_{i}),i=1,\ldots,n$)
$d_{D,2}(\mathbf{h},\hat{\mathbf{h}})=\mathbb{E}_{\sigma_{i_{1}},i_{1}=1,...,n}|\frac{1}{C_{n}^{4}}\sum_{1\leq i_{1}<i_{2}<i_{3}<i_{4}\leq n}\sigma_{i_{1}}(k_{\mathbf{h}}-k_{\hat{\mathbf{h}}})(\tilde{D}_{i_{1}},\ldots,\tilde{D}_{i_{4}})|,$
where $\sigma_{i_{1}}\sim Rad(\frac{1}{2})$ i.i.d and $k_{\mathbf{h}}(\tilde{D}_{i_{1}},\ldots,\tilde{D}_{i_{4}})=k((\mathbf{h}(\mathbf{X}_{i_{1}}),S_{i_{1}}),\ldots,(\mathbf{h}(\mathbf{X}_{i_{4}}),S_{i_{4}}))$.
Condition on $\tilde{D}_{i},i=1,\ldots,n$, let $\mathfrak{C}(\mathcal{H},d_{D,2},\delta))$ be the covering number Giné and Nickl [78, page 41] of $\mathcal{H}$ with respect to the empirical distance $d_{D,2}$ at scale of $\delta>0$. Denote $\mathcal{H}_{\delta}$ as the covering set of $\mathcal{H}$ with with cardinality of $\mathfrak{C}(\mathcal{H},d_{D,2},\delta))$.
Lemma 4.
If $\xi_{i},i=1,...m$ are $m$ finite linear combinations of Rademacher variables $\epsilon_{j},j=1,..J$. There exists some $C_{0},C_{1},C_{2}>0$, Then
$\mathbb{E}_{\epsilon_{j},j=1,...J}\max_{1\leq i\leq m}|\xi_{i}|\leq C_{2}(\log m)^{1/2}\max_{1\leq i\leq m}\left(\mathbb{E}\xi_{i}^{2}\right)^{1/2}.$ (20)
Proof.
By De la Pena and Giné [79][Corollary 3.2.6], we know that $\xi_{i}$ is all belongs to $\mathcal{L}_{\phi_{2}}$ (a space of Orlicz space with $\phi_{2}(x)=exp(x^{2})$) for their corresponding Orlicz norm [79, page 36] is $C_{0}\mathrm{E}[\xi_{i}^{2}]^{\frac{1}{2}}$. So it’s still a Sub-Gaussian r.v with parameter $C_{1}\mathrm{E}[\xi_{i}^{2}]^{\frac{1}{2}}$, and by De la Pena and Giné [79][inequality (4.3.1)] with $\Phi(x)=\exp(\rho x)$ we can get that
$\mathbb{E}_{\epsilon_{j},j=1,...J}\max_{1\leq i\leq m}|\xi_{i}|\leq\frac{\log m}{\rho}+\frac{\log\max_{1\leq i\leq m}\exp{\frac{\rho^{2}C_{1}\mathrm{E}[\xi_{i}^{2}]^{\frac{1}{2}}}{2}}}{\rho}.$
, then we can choose a optimal $\rho$ to get the wanted result. ∎
Lemma 5.
There exists some $C_{3},C_{4},C_{5},C_{6}>0$, making the following inequality satisfied.
$\mathcal{L}_{2}\leq C_{6}\sqrt{r}\sqrt{\frac{{\rm Pdim}(\mathcal{H}_{1})}{n}},$ (21)
where $\mathcal{H}_{1}$ means the neural network function class $\mathcal{N}\mathcal{N}_{d,1}(W_{1},L_{1},K_{1})$ which can be seen as the projection to one-dimensional of $\mathcal{H}$.
Proof.
Base the definition of $\mathfrak{C}(\mathcal{H},d_{D,2},\delta))$ and $\mathcal{H}_{\delta}$,
$\mathcal{L}_{2}=$ $\mathbb{E}\mathbb{E}_{\sigma_{i_{1}}}[\sup_{h\in\mathcal{H}}|\frac{1}{C_{n}^{4}}\sum_{1\leq i_{1}<i_{2}<i_{3}<i_{4}\leq n}\sigma_{i_{1}}\bar{k}_{\mathbf{h}}(\tilde{D}_{i_{1}},\tilde{D}_{i_{2}},\tilde{D}_{i_{3}},\tilde{D}_{i_{4}})|]$
$\leq\inf_{0<\delta<1}\{\delta+\mathbb{E}\mathbb{E}_{\sigma_{i_{1}}}[\sup_{\mathbf{h}\in\mathcal{H}_{\delta}}|\frac{1}{C_{n}^{4}}\sum_{1\leq i_{1}<i_{2}<i_{3}<i_{4}\leq n}\sigma_{i_{1}}\bar{k}_{\mathbf{h}}(\tilde{D}_{i_{1}},\tilde{D}_{i_{2}},\tilde{D}_{i_{3}},\tilde{D}_{i_{4}})|]\}$
$\leq\inf_{0<\delta<1}\{\delta+C_{3}\frac{1}{C_{n}^{4}}\mathbb{E}[(\log\mathfrak{C}(\mathcal{H},d_{D,2},\delta))^{1/2}\max_{\mathbf{h}\in\mathcal{H}_{\delta}}[\sum_{i_{1}=1}^{n}\sum_{i_{2}<i_{3}<i_{4}}(\bar{k}_{\mathbf{h}}(\tilde{D}_{i_{1}},\tilde{D}_{i_{2}},\tilde{D}_{i_{3}},\tilde{D}_{i_{4}}))^{2}]^{1/2}]\}$
$\leq\inf_{0<\delta<1}\{\delta+C_{3}C_{4}\mathbb{E}[(\log\mathfrak{C}(\mathcal{H},d_{D,2},\delta))^{1/2}\frac{1}{C_{n}^{4}}[\frac{n(n!)^{2}}{((n-3)!)^{2}}]^{1/2}]\}$
$\leq\inf_{0<\delta<1}\{\delta+2C_{3}C_{4}\mathbb{E}[(\log\mathfrak{C}(\mathcal{H},d_{D,2},\delta))^{1/2}/\sqrt{n}]\}$
$\leq\inf_{0<\delta<1}\{\delta+2C_{3}C_{4}\mathbb{E}[(\log\mathfrak{C}(\mathcal{H},C_{5}d_{D,\infty},\delta))^{1/2}/\sqrt{n}]\}$
$\leq\inf_{0<\delta<1}\{\delta+2C_{3}C_{4}\mathbb{E}[(\log\mathfrak{C}(\mathcal{H},d_{D,\infty},\delta/C_{5}))^{1/2}/\sqrt{n}]\}$
$\leq\inf_{0<\delta<1}\{\delta+2C_{3}C_{4}\sqrt{r}\mathbb{E}[(\log\mathfrak{C}(\mathcal{H}_{1},d_{D,\infty},\delta/C_{5}))^{1/2}/\sqrt{n}]\}$
$\leq\inf_{0<\delta<1}\{\delta+2C_{3}C_{4}\sqrt{r}({\rm Pdim}(\mathcal{H}_{1})\log(\frac{2C_{5}en}{\delta{\rm Pdim}(\mathcal{H}_{1})}))^{1/2}/\sqrt{n}\}$
$\leq C_{6}\sqrt{r}\sqrt{\frac{{\rm Pdim}(\mathcal{H}_{1})}{n}},$
where $d_{D,\infty}(\mathbf{h}_{1},\mathbf{h}_{2})=\sum_{i=1}^{n}\frac{1}{n}\|\mathbf{h}_{1}(\mathbf{x}_{i})-\mathbf{h}_{2}(\mathbf{x}_{2})\|_{\infty}$. The proof is base on the proof on Huang et al. [75, Lemma 8.7], and the second inequality is base on the lemma 4.
As for the third last inequality, when estimating the covering number of the family of neural network functions, it is commonly approached by employing concepts such as pseudo dimension or VC dimension. Denote the pseudo dimension of $\mathcal{F}$ by ${\rm Pdim}(\mathcal{F})$, by Anthony et al. [80, Theorem 12.2], for $n\geq{\rm Pdim}(\mathcal{F})$,
$\mathfrak{C}(\mathcal{F},d_{D,\infty},\delta)\leq\Big{(}\frac{2eB_{\mathcal{F}}n}{{\rm Pdim}(\mathcal{F})\delta}\Big{)}^{{\rm Pdim}(\mathcal{F})}.$
Here $B_{\mathcal{F}}$ is the upper bound of funtions in $\mathcal{F}$. However,we can establish its validity regardless of the relative magnitudes of $n$ and ${\rm Pdim}(\mathcal{B}_{pr}(B)\circ\mathcal{H})$, as demonstrated by the proof of Huang et al. [81, Corollary 35].
For the last inequality we set the $\delta=C_{4}\sqrt{\frac{{\rm Pdim}(\mathcal{H}_{1})}{n}}<1$. ∎
Lemma 6.
$\mathcal{L}_{3}$ $=\mathbb{E}[\sup_{\mathbf{h}\in\mathcal{H}}|\hat{\mathcal{W}_{1}}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})-\mathcal{W}_{1}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})|]$ (22)
$\leq 2\sup_{h\in\mathcal{H}^{1}}\inf_{g\in\mathcal{G}}\|g-h\|+2\mathbb{E}[\sup_{\mathbf{h}\in\mathcal{H},g\in\mathcal{G}}|\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}g(\mathbf{h}(\mathbf{X}_{i}))|]$
$\quad+2\mathbb{E}[\sup_{g\in\mathcal{G}}|\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}g(\xi_{i})|].$
Proof.
Let $\overline{\mathcal{W}_{1}}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r}):=\sup_{g\in\mathcal{G}_{1}}\frac{1}{K_{2}}\mathbb{E}_{\mathbf{x}\sim\mathbf{h}(\mathbf{X})}[g(\mathbf{X})]-\mathbb{E}_{\mathbf{s}\sim\mathcal{U}_{r}}[g(z)]$ be the population form of $\hat{\mathcal{W}_{1}}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})$, then we can derive the proof as followed:
$\mathcal{L}_{3}$ $=\mathbb{E}[\sup_{\mathbf{h}\in\mathcal{H}}|\hat{\mathcal{W}_{1}}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})-\mathcal{W}_{1}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})|]$
$\leq\mathbb{E}[\sup_{\mathbf{h}\in\mathcal{H}}|\hat{\mathcal{W}_{1}}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})-\overline{\mathcal{W}_{1}}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})+\overline{\mathcal{W}_{1}}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})-\mathcal{W}_{1}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})|]$
$\leq\frac{2}{K_{2}}\sup_{h\in\mathcal{H}^{1}}\inf_{g\in\mathcal{G}}\|g-h\|+\frac{1}{K_{2}}\mathbb{E}[\sup_{h\in\mathcal{H}^{1}}|\sup_{g\in\mathcal{G}}\frac{1}{n}\sum_{i=1}^{n}g(\mathbf{h}(\mathbf{X}_{i}))-\frac{1}{n}\sum_{i=1}^{n}g(\xi_{i})$
$\quad-\sup_{g\in\mathcal{G}}\mathbb{E}_{\mathbf{x}\sim\mathbf{h}(\mathbf{X})}[g(\mathbf{x})]-\mathbb{E}_{\mathbf{z}\sim\mathcal{U}_{r}}[g(\mathbf{z})]|]$
$\leq\frac{2}{K_{2}}\sup_{h\in\mathcal{H}^{1}}\inf_{g\in\mathcal{G}}\|g-h\|+\frac{1}{K_{2}}\mathbb{E}[\sup_{h\in\mathcal{H}^{1}}\sup_{g\in\mathcal{G}}|\mathbb{E}_{\mathbf{x}\sim\mathbf{h}(\mathbf{X})}[g(\mathbf{x})]-\frac{1}{n}\sum_{i=1}^{n}g(\mathbf{h}(\mathbf{X}_{i}))$
$\quad+\frac{1}{n}\sum_{i=1}^{n}g(\xi_{i})-\mathbb{E}_{\mathbf{z}\sim\mathcal{U}_{r}}[g(\mathbf{z})]|]$
$\leq\frac{2}{K_{2}}\sup_{h\in\mathcal{H}^{1}}\inf_{g\in\mathcal{G}}\|g-h\|+\frac{1}{K_{2}}\mathbb{E}[\sup_{h\in\mathcal{H}^{1}}\sup_{g\in\mathcal{G}}|\mathbb{E}_{\mathbf{x}\sim\mathbf{h}(\mathbf{X})}[g(\mathbf{x})]-\frac{1}{n}\sum_{i=1}^{n}g(\mathbf{h}(\mathbf{X}_{i}))$
$\quad+|\frac{1}{n}\sum_{i=1}^{n}g(\xi_{i})-\mathbb{E}_{\mathbf{z}\sim\mathcal{U}_{r}}[g(\mathbf{z})]|]$
$\leq\frac{2}{K_{2}}\sup_{h\in\mathcal{H}^{1}}\inf_{g\in\mathcal{G}}\|g-h\|+\frac{1}{K_{2}}\mathbb{E}[\sup_{\mathbf{h}\in\mathcal{H},g\in\mathcal{G}}|\mathbb{E}_{\mathbf{x}\sim\mathbf{h}(\mathbf{X})}[g(\mathbf{x})]-\frac{1}{n}\sum_{i=1}^{n}g(\mathbf{h}(\mathbf{X}_{i}))|$
$\quad+\sup_{g\in\mathcal{G}}|\mathbb{E}_{\mathbf{z}\sim\mathcal{U}_{r}}[g(\mathbf{z})]-\frac{1}{n}\sum_{i=1}^{n}g(\xi_{i})|]$
$\leq\frac{2}{K_{2}}\sup_{h\in\mathcal{H}^{1}}\inf_{g\in\mathcal{G}}\|g-h\|+\frac{2}{K_{2}}\mathbb{E}[\sup_{\mathbf{h}\in\mathcal{H},g\in\mathcal{G}}|\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}g(\mathbf{h}(\mathbf{X}_{i}))|]$
$\quad+\frac{2}{K_{2}}\mathbb{E}[\sup_{g\in\mathcal{G}}|\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}g(\xi_{i})|].$
The first inequality is due to the metric property. The second inequality bounding the term $\overline{\mathcal{W}_{1}}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})-\mathcal{W}_{1}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})$ by Huang et al. [81, lemma 24]. The third inequality is due to the $|\sup_{a}f-\sup_{a}g|\leq\sup_{a}|f-g|$.The second last and last inequality is again by applying chain and symmetrization technique as proof of lemma 3. ∎
We now present the proof of our main result for upstream.
Proof of Theorem 1.
Combine all the lemma proofed in Section B.1.1 and B.1.2, we can get that
$\mathbb{E}[R(\hat{\mathbf{F}},\hat{\mathbf{h}})-R(\mathbf{F}^{*},\mathbf{h}^{*})]$ $\leq\frac{2B^{2}}{\sqrt{n}}+(4B^{2}+32p+1)\|\mathbf{h}-\mathbf{h}^{*}\|+\frac{2}{K_{2}}\sup_{f\in\mathcal{H}^{1}}\inf_{g\in\mathcal{G}}\|g-f\|$ (23)
$\quad+8B\mathbb{E}[\sup_{(\mathbf{F},\mathbf{h})\in\mathcal{B}_{pr}(B)\circ\mathcal{H}}|\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\mathbf{F}_{S_{i}}\mathbf{h}(\mathbf{X}_{S_{i},i})|]+C_{6}\sqrt{r}\sqrt{\frac{{\rm Pdim}(\mathcal{H}_{1})}{n}}$
$\quad+\frac{2}{K_{2}}\mathbb{E}[\sup_{\mathbf{h}\in\mathcal{H},g\in\mathcal{G}}|\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}g(\mathbf{h}(\mathbf{X}_{i}))|]+\frac{2}{K_{2}}\mathbb{E}[\sup_{g\in\mathcal{G}}|\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}g(\xi_{i})|]\}.$
Based on [82], the pseudo dimension (VC dimension) of $\mathcal{F}$ with piecewise-linear activation function can be further contained and represented by its depth $L$ and width $W$, i.e., ${\rm Pdim}(\mathcal{F})=O(W^{2}L^{2}\log(W^{2}L)$.
By ${\rm Pdim}(\mathcal{F})=O(W^{2}L^{2}\log(W^{2}L)$ for $\mathcal{F}=\mathcal{N}\mathcal{N}(W,L)$ and Theorem 5, 6, (23) can be further deduced as followed:
$\mathbb{E}[R(\hat{\mathbf{F}},\hat{\mathbf{h}})-R(\mathbf{F}^{*},\mathbf{h}^{*})]$ $\precsim\frac{1}{\sqrt{n}}+K_{1}^{-\frac{\beta}{d+1}}+K_{2}^{-\frac{\beta}{r+1}}$
$\quad+\frac{K_{1}}{\sqrt{n}}+\sqrt{r}\sqrt{\frac{(W_{1}L_{1})^{2}\log(W_{1}^{2}L_{1})\log n}{n}}$
$\quad+\frac{K_{1}}{\sqrt{n}}+\frac{1}{\sqrt{n}}$
$\precsim K_{1}^{-\frac{\beta}{d+1}}+K_{2}^{-\frac{\beta}{r+1}}+\frac{K_{1}^{\frac{2d+\beta}{2d+2}}}{\sqrt{n}}+\frac{K_{1}}{\sqrt{n}}+\frac{1}{\sqrt{n}}$
$\precsim n^{-\frac{\beta}{2(d+1+\beta)}}\mathbb{I}_{\beta\leq 2}+n^{-\frac{\beta}{2d+3\beta}}\log n\mathbb{I}_{\beta>2}$
$\precsim\tilde{O}\left(n^{-\frac{\beta}{2(d+1+\beta)}}\mathbb{I}_{\beta\leq 2}+n^{-\frac{\beta}{2d+3\beta}}\mathbb{I}_{\beta>2}\right),$
where we set $K_{1}\asymp n^{\frac{d+1}{2(d+1+\beta)}}$, $K_{2}\asymp n^{\frac{r+1}{2(d+1+\beta)}}$, and $W_{1}\asymp K_{1}^{\frac{2d+\beta}{2d+2}}$ when $\beta\leq 2$.
When $\beta>2$, we set $K_{1}\asymp n^{\frac{d+1}{2d+3\beta}}$, $K_{2}\asymp n^{\frac{d+1}{2d+3\beta}}$, and $W_{1}\asymp K_{1}^{\frac{2d+\beta}{2d+2}}$.
B.2 Proof of Theorem 2
We now present the proof of result about sparsity of ERM solutions in upstream. Unlike the proof of Theorem 1 presented earlier, for the sake of explanatory coherence, we have positioned the statements and proofs of the necessary lemmas after the proof of the corresponding theorem in the subsequent demonstration.
Proof of Theorem 2.
For simplicity, we denote $n^{-t(\beta)}=n^{-\frac{\beta}{2(d+1+\beta)}}\mathbb{I}_{\beta\leq 2}+n^{-\frac{\beta}{2d+3\beta}}\log n\mathbb{I}_{\beta>2}$.
By the definition of $(\hat{\mathbf{F}},\hat{\mathbf{h}})$ from (12), we know that $\hat{R}(\hat{\mathbf{F}},\hat{\mathbf{h}})\leq\hat{R}(\mathbf{f}_{0},\hat{\mathbf{h}})$.
$\sum_{i=1}^{p}\sum_{j=1}^{n_{p}}\frac{1}{n}(Y_{i,j}-\hat{\mathbf{F}}_{i}\hat{\mathbf{h}}(\mathbf{X}_{i,j}))^{2}+\mu\sum_{i=1}^{p}\|\hat{\mathbf{F}}_{i}\|_{1}\leq\sum_{i=1}^{p}\sum_{j=1}^{n_{p}}\frac{1}{n}(Y_{i,j}-\mathbf{F}^{*}_{i}\hat{\mathbf{h}}(\mathbf{X}_{i,j}))^{2}+\mu\sum_{i=1}^{p}\|\mathbf{F}^{*}_{i}\|_{1},$
$\sum_{i=1}^{p}\sum_{j=1}^{n_{p}}\frac{1}{n}(Y_{i,j}-\hat{\mathbf{F}}_{i}\hat{\mathbf{h}}(\mathbf{X}_{i,j}))^{2}\leq\sum_{i=1}^{p}\sum_{j=1}^{n_{p}}\frac{1}{n}(Y_{i,j}-\mathbf{F}^{*}_{i}\hat{\mathbf{h}}(\mathbf{X}_{i,j}))^{2}+\mu\sum_{i=1}^{p}\|\hat{\mathbf{F}}_{i}-\mathbf{F}^{*}_{i}\|_{1}.$ (24)
By the equation
$\sum_{i=1}^{p}\sum_{j=1}^{n_{p}}\frac{1}{n}(Y_{i,j}-\hat{\mathbf{F}}_{i}\hat{\mathbf{h}}(\mathbf{X}_{i,j}))^{2}$ $=\sum_{i=1}^{p}\sum_{j=1}^{n_{p}}\frac{1}{n}(Y_{i,j}-\mathbf{F}^{*}_{i}\hat{\mathbf{h}}(\mathbf{X}_{i,j})+\mathbf{F}^{*}_{i}\hat{\mathbf{h}}(\mathbf{X}_{i,j})-\hat{\mathbf{F}}_{i}\hat{\mathbf{h}}(\mathbf{X}_{i,j}))^{2}$ (25)
$=\sum_{i=1}^{p}\sum_{j=1}^{n_{p}}\frac{1}{n}(Y_{i,j}-\mathbf{F}^{*}_{i}\hat{\mathbf{h}}(\mathbf{X}_{i,j}))^{2}$
$+\sum_{i=1}^{p}\sum_{j=1}^{n_{p}}\frac{2}{n}(Y_{i,j}-\mathbf{F}^{*}_{i}\hat{\mathbf{h}}(\mathbf{X}_{i,j}))(\mathbf{F}^{*}_{i}\hat{\mathbf{h}}(\mathbf{X}_{i,j})-\hat{\mathbf{F}}_{i}\hat{\mathbf{h}}(\mathbf{X}_{i,j}))$
$+\sum_{i=1}^{p}\sum_{j=1}^{n_{p}}\frac{1}{n}((\mathbf{F}^{*}_{i}-\hat{\mathbf{F}}_{i})\hat{\mathbf{h}}(\mathbf{X}_{i,j})^{2},$
after estimating the last two term in right part of upper equation. We can replace the left part in (24) by (25) to get the desired result by selecting proper $\mu$.
Due to the lemma 8 below, by selecting $\mu=\frac{\gamma}{2Bpr}n^{-t(\beta)}$, the mid term of (25) can be bounded as followed with probability at least $1-2\delta$ over the draw of dataset $\mathcal{D}$ (w.h.p):
$\sum_{i=1}^{p}\sum_{j=1}^{n_{p}}\frac{2}{n}(Y_{i,j}-\mathbf{F}^{*}_{i}\hat{\mathbf{h}}(\mathbf{X}_{i,j}))(\mathbf{F}^{*}_{i}\hat{\mathbf{h}}(\mathbf{X}_{i,j})-\hat{\mathbf{F}}_{i}\hat{\mathbf{h}}(\mathbf{X}_{i,j}))$
$=$ $\sum_{i=1}^{p}\sum_{j=1}^{n_{p}}\frac{2}{n}(\mathbf{F}^{*}_{i}\mathbf{h}^{*}(\mathbf{X}_{i,j})-\mathbf{F}^{*}_{i}\hat{\mathbf{h}}(\mathbf{X}_{i,j})+\epsilon_{i,j})(\mathbf{F}^{*}_{i}-\hat{\mathbf{F}}_{i})\hat{\mathbf{h}}(\mathbf{X}_{i,j})$
$=$ $\sum_{i=1}^{p}\sum_{j=1}^{n_{p}}\frac{2}{n}(\mathbf{F}^{*}_{i}(\mathbf{h}^{*}(\mathbf{X}_{i,j})-\hat{\mathbf{h}}(\mathbf{X}_{i,j}))+\epsilon_{i,j})(\mathbf{F}^{*}_{i}-\hat{\mathbf{F}}_{i})\hat{\mathbf{h}}(\mathbf{X}_{i,j})$
$\leq$ $\sum_{i=1}^{p}\|\mathbf{F}^{*}_{i}-\hat{\mathbf{F}}_{i}\|_{1}\|\sum_{j=1}^{n_{p}}\frac{2}{n}\mathbf{F}^{*}_{i}(\mathbf{h}^{*}(\mathbf{X}_{i,j})-\hat{\mathbf{h}}(\mathbf{X}_{i,j}))\hat{\mathbf{h}}(\mathbf{X}_{i,j})\|_{\infty}$
$+$ $\sum_{i=1}^{p}\|\mathbf{F}^{*}_{i}-\hat{\mathbf{F}}_{i}\|_{1}\|\sum_{j=1}^{n_{i}}\frac{2}{n}\epsilon_{i,j}\hat{\mathbf{h}}(\mathbf{X}_{i,j})\|_{\infty}$
$\overset{w.h.p}{\leq}$ $\sum_{i=1}^{p}\|\mathbf{F}^{*}_{i}-\hat{\mathbf{F}}_{i}\|_{1}\left((1+2B)(1+2B_{2})\gamma n^{-t(\beta)}+(2B+1)\sqrt{\frac{\log\frac{1}{\delta}}{2n}}+2\sqrt{\frac{2\log\frac{2r}{\delta}}{n_{i}}}\right),$
where $\epsilon_{i,j}$ here is the i.i.d sample of $\varepsilon_{1}$ and for simplicity we omit the footnote from $\epsilon_{1,i,j}$ to $\epsilon_{i,j}$.
By A(4) in the Assumption 2, we can get
$\frac{1}{n}\sum_{i=1}^{p}\sum_{j=1}^{n_{p}}((\mathbf{F}^{*}_{i}-\hat{\mathbf{F}}_{i})\hat{\mathbf{h}}(\mathbf{X}_{i,j}))^{2}=\sum_{i=1}^{p}\|(\mathbf{F}^{*}_{i}-\hat{\mathbf{F}}_{i})\hat{\mathbf{H}}_{i}\|_{2}^{2}\geq\sum_{i=1}^{p}B_{1}\|\mathbf{F}^{*}_{i}-\hat{\mathbf{F}}_{i}\|_{2}^{2}.$
So the (24) can be further deduced as followed:
$\sum_{i=1}^{p}B_{1}\|\mathbf{F}^{*}_{i}-\hat{\mathbf{F}}_{i}\|_{2}^{2}\overset{w.h.p}{\leq}\mu\sum_{i=1}^{p}\|\hat{\mathbf{F}}_{i}-\mathbf{F}^{*}_{i}\|_{1}$
$\quad+\sum_{i=1}^{p}\|\mathbf{F}^{*}_{i}-\hat{\mathbf{F}}_{i}\|_{1}\left((1+2B)(1+2B_{2})\gamma n^{-t(\beta)}+(2B+1)\sqrt{\frac{\log\frac{1}{\delta}}{2n}}+2\sqrt{\frac{2\log\frac{2r}{\delta}}{n_{i}}}\right),$
By replacing $(2B+1)\sqrt{\frac{\log\frac{1}{\delta}}{2n}}+2\sqrt{\frac{2\log\frac{2r}{\delta}}{n_{i}}}$ by $\mathcal{O}\left(\frac{1}{\sqrt{n}}\right)$ for simplicity
$~{}~{}~{}~{}\sum_{i=1}^{p}$ $B_{1}\|\mathbf{F}^{*}_{i}-\hat{\mathbf{F}}_{i}\|_{2}^{2}$
$\overset{w.h.p}{\leq}$ $\sum_{i=1}^{p}\Big{(}\big{(}(1+2B)(1+2B_{2})+\frac{1}{2Bpr}\big{)}\gamma n^{-t(\beta)}+\mathcal{O}\big{(}\frac{1}{\sqrt{n}}\big{)}\Big{)}\|\hat{\mathbf{F}}_{i}-\mathbf{F}^{*}_{i}\|_{1}$
$\overset{w.h.p}{\leq}$ $\sum_{i=1}^{p}\sqrt{r}\Big{(}\big{(}(1+2B)(1+2B_{2})+\frac{1}{2Bpr}\big{)}\gamma n^{-t(\beta)}+\mathcal{O}\big{(}\frac{1}{\sqrt{n}}\big{)}\Big{)}\|\hat{\mathbf{F}}_{i}-\mathbf{F}^{*}_{i}\|_{2}.$
Then we have
$B_{1}\sqrt{\sum_{i=1}^{p}\|\mathbf{F}^{*}_{i}-\hat{\mathbf{F}}_{i}\|_{2}^{2}}\overset{w.h.p}{\leq}\sum_{i=1}^{p}\sqrt{rp}\Big{(}\big{(}(1+2B)(1+2B_{2})+\frac{1}{2Bpr}\big{)}\gamma n^{-t(\beta)}+\mathcal{O}\big{(}\frac{1}{\sqrt{n}}\big{)}\Big{)}.$
Thus we get the wanted result which means w.p.t that the sparsity of $\hat{\mathbf{F}}_{i}$ is similar with $\mathbf{F}_{i}^{*}$. ∎
First, we state the following lemma, which is used in the proof of Lemma 8.
Lemma 7.
$\mathbb{E}[\mathcal{W}_{1}(\hat{\mathbf{h}}(\mathbf{X}^{{}^{\prime}}),\mathbf{h}^{*}(\mathbf{X}^{{}^{\prime}}))+\mu\sum_{i=1}^{p}(\|\hat{\mathbf{F}}_{i}\|_{1}-\|\mathbf{F}^{*}_{i}\|_{1})]\leq\gamma n^{-t(\beta)},$
where $\gamma>0$ is a constant.
Proof of Lemma 7.
The Theorem 1 shows $\mathbb{E}[R(\hat{\mathbf{F}},\hat{\mathbf{h}})-R(\mathbf{F}^{*},\mathbf{h}^{*})]\precsim n^{-t(\beta)}$ and by the definition of $\mathbf{h}^{*}$, we know that
$n^{-t(\beta)}\gtrsim$ $\mathbb{E}[(Y^{{}^{\prime}}_{S}-\hat{\mathbf{F}_{S}}\hat{\mathbf{h}}(\mathbf{X}^{{}^{\prime}}_{S}))^{2}+\mathcal{V}[\hat{\mathbf{h}}(\mathbf{X}^{{}^{\prime}}),S]+\mathcal{W}_{1}(\hat{\mathbf{h}}(\mathbf{X}^{{}^{\prime}}),\mathcal{U}_{r})-(Y^{{}^{\prime}}_{S}-\mathbf{F}^{*}_{S}\mathbf{h}^{*}(\mathbf{X}^{{}^{\prime}}_{S}))^{2}$
$+\mu\sum_{i=1}^{p}(\|\hat{\mathbf{F}}_{i}\|_{1}-\|\mathbf{F}^{*}_{i}\|_{1})]$
$\gtrsim$ $\mathbb{E}[(\varepsilon^{{}^{\prime}}_{1}+\mathbf{F}^{*}_{S}\mathbf{h}^{*}(\mathbf{X}^{{}^{\prime}}_{S})-\hat{\mathbf{F}_{S}}\hat{\mathbf{h}}(\mathbf{X}^{{}^{\prime}}_{S}))^{2}+\mathcal{W}_{1}(\hat{\mathbf{h}}(\mathbf{X}^{{}^{\prime}}),\mathbf{h}^{*}(\mathbf{X}^{{}^{\prime}}))-(\varepsilon^{{}^{\prime}}_{1})^{2}$
$+\mu\sum_{i=1}^{p}(\|\hat{\mathbf{F}}_{i}\|_{1}-\|\mathbf{F}^{*}_{i}\|_{1})]$
$=$ $\mathbb{E}[2\varepsilon^{{}^{\prime}}_{1}(\mathbf{F}^{*}_{S}\mathbf{h}^{*}(\mathbf{X}^{{}^{\prime}}_{S})-\hat{\mathbf{F}_{S}}\hat{\mathbf{h}}(\mathbf{X}^{{}^{\prime}}_{S}))+\mathcal{W}_{1}(\hat{\mathbf{h}}(\mathbf{X}^{{}^{\prime}}),\mathbf{h}^{*}(\mathbf{X}^{{}^{\prime}}))+\mu\sum_{i=1}^{p}(\|\hat{\mathbf{F}}_{i}\|_{1}-\|\mathbf{F}^{*}_{i}\|_{1})]$
$=$ $\mathbb{E}[\mathcal{W}_{1}(\hat{\mathbf{h}}(\mathbf{X}^{{}^{\prime}}),\mathbf{h}^{*}(\mathbf{X}^{{}^{\prime}}))+\mu\sum_{i=1}^{p}(\|\hat{\mathbf{F}}_{i}\|_{1}-\|\mathbf{F}^{*}_{i}\|_{1})]$
Here ${\mathbf{X}^{{}^{\prime}}_{S},Y^{{}^{\prime}}_{S},\varepsilon^{{}^{\prime}}_{1}}$ is the i.i.d copy of ${\mathbf{X}_{S},Y_{S},\varepsilon_{1}}$. For the simplicity of calculation, we can set a constant $\gamma>0$ big enough that $\mathbb{E}[\mathcal{W}_{1}(\hat{\mathbf{h}}(\mathbf{X}^{{}^{\prime}}),\mathbf{h}^{*}(\mathbf{X}^{{}^{\prime}}))+\mu\sum_{i=1}^{p}(\|\hat{\mathbf{F}}_{i}\|_{1}-\|\mathbf{F}^{*}_{i}\|_{1})]\leq\gamma n^{-t(\beta)}$
Lemma 8.
For any $\delta>0$, $i\in[p]$, with probability at least $1-2\delta$ we get:
$\|\sum_{j=1}^{n_{i}}\frac{1}{n}\mathbf{F}^{*}_{i}(\mathbf{h}^{*}(\mathbf{X}_{i,j})-\hat{\mathbf{h}}(\mathbf{X}_{i,j}))\hat{\mathbf{h}}(\mathbf{X}_{i,j})\|_{\infty}\leq$ $B(1+2B_{2})\gamma n^{-t(\beta)}+B\sqrt{\frac{\log\frac{1}{\delta}}{2n}},$
$\|\sum_{j=1}^{n_{i}}\frac{1}{n}\epsilon_{i,j}\hat{\mathbf{h}}(\mathbf{X}_{i,j})\|_{\infty}\leq$ $(1+2B_{2})\gamma n^{-t(\beta)}+\sqrt{\frac{\log\frac{1}{\delta}}{2n}}+\sqrt{\frac{2\log\frac{2r}{\delta}}{n_{i}}},$
where $\gamma$ is a constant big enough.
Proof of Lemma 8.
By A(1) in the Assumption 1, we know that,
$\|\sum_{j=1}^{n_{i}}\frac{1}{n}\mathbf{F}^{*}_{i}(\mathbf{h}^{*}(\mathbf{X}_{i,j})-\hat{\mathbf{h}}(\mathbf{X}_{i,j}))\hat{\mathbf{h}}(\mathbf{X}_{i,j})\|_{\infty}\leq$ $B\sum_{j=1}^{n_{i}}\frac{1}{n}\|(\mathbf{h}^{*}(\mathbf{X}_{i,j})-\hat{\mathbf{h}}(\mathbf{X}_{i,j}))\|_{1}$
$\|\sum_{j=1}^{n_{i}}\frac{1}{n}\mathbf{F}^{*}_{i}(\mathbf{h}^{*}(\mathbf{X}_{i,j})-\hat{\mathbf{h}}(\mathbf{X}_{i,j}))\hat{\mathbf{h}}(\mathbf{X}_{i,j})\|_{\infty}\leq$ $B\sum_{i=1}^{p}\sum_{j=1}^{n_{i}}\frac{1}{n}\|(\mathbf{h}^{*}(\mathbf{X}_{i,j})-\hat{\mathbf{h}}(\mathbf{X}_{i,j}))\|_{1},$
$\|\sum_{j=1}^{n_{i}}\frac{1}{n}\epsilon_{i,j}\hat{\mathbf{h}}(\mathbf{X}_{i,j})\|_{\infty}\leq$ $\sum_{i=1}^{p}\sum_{j=1}^{n_{i}}\frac{1}{n}\|\epsilon_{i,j}\|_{\infty}\|\mathbf{h}^{*}(\mathbf{X}_{i,j})-\hat{\mathbf{h}}(\mathbf{X}_{i,j})\|_{1}$
$+\|\sum_{j=1}^{n_{i}}\frac{1}{n}\epsilon_{i,j}\mathbf{h}^{*}(\mathbf{X}_{i,j})\|_{\infty},$
By lemma 7, we can select $\mu=\frac{\gamma}{2Bpr}n^{-t(\beta)}$ to get that $\mathbb{E}[\mathcal{W}_{1}(\hat{\mathbf{h}}(\mathbf{X}^{{}^{\prime}}),\mathbf{h}^{*}(\mathbf{X}^{{}^{\prime}}))]\leq 2\gamma n^{-t(\beta)}$.
Due to the Theorem 3.3 in [74] and the Theorem 6. For any $\delta>0$, with probability at least $1-\delta$:
$\sum_{i=1}^{p}\sum_{j=1}^{n_{i}}\frac{1}{n}\|\mathbf{h}^{*}(\mathbf{X}_{i,j})-\hat{\mathbf{h}}(\mathbf{X}_{i,j})\|_{1}-\mathbb{E}[\|\mathbf{h}^{*}(\mathbf{X}^{{}^{\prime}})-\hat{\mathbf{h}}(\mathbf{X}^{{}^{\prime}})\|_{1}]$
$\leq$ $2\mathbb{E}[\sup_{\mathbf{h}\in\mathcal{H}}\frac{1}{n}\sum_{i=1}^{p}\sum_{j=1}^{n_{i}}\sigma_{i}\|\mathbf{h}(\mathbf{X}_{i,j})-\mathbf{h}^{*}(\mathbf{X}_{i,j})\|_{1}]+\sqrt{\frac{\log\frac{1}{\delta}}{2n}}$
$\leq$ $2\mathbb{E}[\sup_{\mathbf{h}\in\mathcal{H}}\frac{1}{n}\sum_{i=1}^{p}\sum_{j=1}^{n_{i}}\sigma_{i,j}\|\mathbf{h}(\mathbf{X}_{i,j})-\mathbf{h}^{*}(\mathbf{X}_{i,j})\|_{1}]+\sqrt{\frac{\log\frac{1}{\delta}}{2n}}$
$\leq$ $\sqrt{\frac{\log\frac{1}{\delta}}{2n}}+\frac{2rK_{1}\sqrt{2(L_{1}+2+\log(d+1))}}{\sqrt{n}}$
$\leq$ $\gamma n^{-t(\beta)}+\sqrt{\frac{\log\frac{1}{\delta}}{2n}}.$
The last inequality is due to properly selected $W_{1},L_{1},K_{1}$ during proof of Theorem 1. Then, by Hoeffding’s inequality:
$\mathbb{P}[\|\sum_{j=1}^{n_{i}}\frac{1}{n_{i}}\epsilon_{i,j}\mathbf{h}^{*}(\mathbf{X}_{i,j})\|_{\infty}\geq\delta_{0}]$
$=$ $\mathbb{P}[\|\sum_{j=1}^{n_{i}}\frac{1}{n_{i}}\epsilon_{i,j}\mathbf{h}^{*}(\mathbf{X}_{i,j})-\mathbb{E}[\varepsilon_{1}\mathbf{h}^{*}(\mathbf{X})]\|_{\infty}\geq\delta_{0}]$
$=$ $\mathbb{P}[\exists k\in[r],|\sum_{j=1}^{n_{i}}\frac{1}{n_{i}}\epsilon_{i,j}\mathbf{h}^{*}_{k}(\mathbf{X}_{i,j})|\geq\delta_{0}]$
$=$ $\mathbb{P}[\exists k\in[r],|\sum_{j=1}^{n_{i}}\frac{1}{n_{i}}\epsilon_{i,j}\mathbf{h}^{*}_{k}(\mathbf{X}_{i,j})-\mathbb{E}[\varepsilon_{1}\mathbf{h}^{*}_{k}(X)]|\geq\delta_{0}]$
$\leq$ $2r\exp(-\frac{\delta_{0}^{2}n_{i}}{2}).$
Then we can set the $\delta_{0}=\sqrt{\frac{2\log\frac{2r}{\delta}}{n_{i}}}$ here, then with probability at least $1-\delta$, we know $\|\sum_{j=1}^{n_{i}}\frac{1}{n_{i}}\epsilon_{i,j}\mathbf{h}^{*}(\mathbf{X}_{i,j})\|_{\infty}\leq\sqrt{\frac{2\log\frac{2r}{\delta}}{n_{i}}}$.
The by A(5) in the Assumption 2, with probability at least $1-2\delta$ we get:
$\|\sum_{j=1}^{n_{i}}\frac{1}{n}\mathbf{F}^{*}_{i}(\mathbf{h}^{*}(\mathbf{X}_{i,j})-\hat{\mathbf{h}}(\mathbf{X}_{i,j}))\hat{\mathbf{h}}(\mathbf{X}_{i,j})\|_{\infty}$
$\leq$ $B\sum_{j=1}^{n_{i}}\frac{1}{n}\|\mathbf{h}^{*}(\mathbf{X}_{i,j})-\hat{\mathbf{h}}(\mathbf{X}_{i,j})\|_{1}-B\mathbb{E}[\|\mathbf{h}^{*}(\mathbf{X}^{{}^{\prime}})-\hat{\mathbf{h}}(\mathbf{X}^{{}^{\prime}})\|_{1}]+B\mathbb{E}[\|\mathbf{h}^{*}(\mathbf{X}^{{}^{\prime}})-\hat{\mathbf{h}}(\mathbf{X}^{{}^{\prime}})\|_{1}]$
$\leq$ $B\sum_{j=1}^{n_{i}}\frac{1}{n}\|\mathbf{h}^{*}(\mathbf{X}_{i,j})-\hat{\mathbf{h}}(\mathbf{X}_{i,j})\|_{1}-B\mathbb{E}[\|\mathbf{h}^{*}(\mathbf{X}^{{}^{\prime}})-\hat{\mathbf{h}}(\mathbf{X}^{{}^{\prime}})\|_{1}]+BB_{2}\mathbb{E}[\mathcal{W}_{1}(\hat{\mathbf{h}}(\mathbf{X}^{{}^{\prime}}),\mathbf{h}^{*}(\mathbf{X}^{{}^{\prime}}))]$
$\leq$ $B(1+2B_{2})\gamma n^{-t(\beta)}+B\sqrt{\frac{\log\frac{1}{\delta}}{2n}}.$
And following the same way before
$\|\sum_{j=1}^{n_{i}}\frac{1}{n}\epsilon_{i,j}\hat{\mathbf{h}}(\mathbf{X}_{i,j})\|_{\infty}\leq$ $\sum_{i=1}^{p}\sum_{j=1}^{n_{i}}\frac{1}{n}\|\epsilon_{i,j}\|_{\infty}\|\mathbf{h}^{*}(\mathbf{X}_{i,j})-\hat{\mathbf{h}}(\mathbf{X}_{i,j})\|_{1}+\|\sum_{j=1}^{n_{i}}\frac{1}{n}\epsilon_{i,j}\mathbf{h}^{*}(\mathbf{X}_{i,j})\|_{\infty}$
$\leq$ $(1+2B_{2})\gamma n^{-t(\beta)}+\sqrt{\frac{\log\frac{1}{\delta}}{2n}}+\sqrt{\frac{2\log\frac{2r}{\delta}}{n_{i}}}.$
Appendix C Proofs of Section 3.2
Next, we list the proof in the fine-tuning, which mainly uses the special property of training loss to induce the control of the excess error of the downstream task by upstream task.
C.1 Proof the Theorem 3
proof the Theorem 3.
In the following proof, for the simplicity of analysis, we incorporated the independence penalty into $\ell(\cdot,\cdot)$ leveraging its $L_{0}$-Lipschitz property. Also, for the simplicity of notion, we denote the function class $\mathcal{Q}\circ\mathcal{B}_{dd^{*}}(B)$ as $\mathcal{Q}_{dd^{*}}:=\{Q_{q,\mathbf{A}}(\cdot)=q(\mathbf{A}\cdot)|q\in\mathcal{Q},\mathbf{A}\in\mathcal{B}_{dd^{*}}(B)\}$. By doing this, we can save much efforts about notation in the analysis without losing it’s coherence.
$\mathbb{E}[\mathcal{L}(\hat{\mathbf{h}},\hat{Q},\hat{\mathbf{F}}_{T})-\mathcal{L}(\mathbf{h}^{*},Q^{*},\mathbf{F}_{T}^{*})]$ (26)
$\leq$ $\mathbb{E}[\mathcal{L}(\hat{\mathbf{h}},\hat{Q},\hat{\mathbf{F}}_{T})-\hat{\mathcal{L}}(\hat{\mathbf{h}},\hat{Q},\hat{\mathbf{F}}_{T})+\hat{\mathcal{L}}(\hat{\mathbf{h}},\hat{Q},\hat{\mathbf{F}}_{T})-\hat{\mathcal{L}}(\hat{\mathbf{h}},Q_{q,\mathbf{A}^{*}},\mathbf{F}_{T}^{*})$
+^$+\hat{\mathcal{L}}(\hat{\mathbf{h}},Q_{q,\mathbf{A}^{*}},\mathbf{F}_{T}^{*})-\mathcal{L}(\hat{\mathbf{h}},Q_{q,\mathbf{A}^{*}},\mathbf{F}_{T}^{*})+\mathcal{L}(\hat{\mathbf{h}},Q_{q,\mathbf{A}^{*}},\mathbf{F}_{T}^{*})-\mathcal{L}(\mathbf{h}^{*},Q_{q,\mathbf{A}^{*}},\mathbf{F}_{T}^{*})$
$+\mathcal{L}(\mathbf{h}^{*},Q_{q,\mathbf{A}^{*}},\mathbf{F}_{T}^{*})-\mathcal{L}(\mathbf{h}^{*},Q^{*},\mathbf{F}_{T}^{*})]$
$\leq$ $\underbrace{\mathbb{E}[\mathcal{L}(\hat{\mathbf{h}},Q_{q,\mathbf{A}^{*}},\mathbf{F}_{T}^{*})-\mathcal{L}(\mathbf{h}^{*},Q_{q,\mathbf{A}^{*}},\mathbf{F}_{T}^{*})]}_{\mathcal{L}_{4}}$
$+\underbrace{2\mathbb{E}[\sup_{(\mathbf{F},Q_{q,\mathbf{A}})\in\mathcal{B}_{r}(B)\circ\mathcal{Q}_{dd^{*}}}cL(\hat{\mathbf{h}},Q_{q,\mathbf{A}},\mathbf{F})-\hat{\mathcal{L}}(\hat{\mathbf{h}},Q_{q,\mathbf{A}},\mathbf{F})|+\inf_{q\in\mathcal{Q}}\|\mathcal{L}(\mathbf{h}^{*},Q_{q,\mathbf{A}^{*}},\mathbf{F}_{T}^{*})-\mathcal{L}(\mathbf{h}^{*},Q^{*},\mathbf{F}_{T}^{*})]}_{\mathcal{L}_{5}}.$
By lemma 9 below, we have:
$\mathcal{L}_{4}\precsim\tilde{O}\left(n^{-\frac{\beta}{2(d+1+\beta)}}\mathbb{I}_{\beta\leq 2}+n^{-\frac{\beta}{2d+3\beta}}\mathbb{I}_{\beta>2}\right).$
By A(6) and A(7) in Assumption 3, Theorem 5, 6 and following the similar process as the proof of lemma 3:
$\mathcal{L}_{5}=$ $2\mathbb{E}[\sup_{(\mathbf{F},Q)\in\mathcal{B}_{r}(B)\circ\mathcal{Q}_{dd^{*}}}|\mathcal{L}(\hat{\mathbf{h}},Q_{q,\mathbf{A}},\mathbf{F})-\hat{\mathcal{L}}(\hat{\mathbf{h}},Q_{q,\mathbf{A}},\mathbf{F})|+\inf_{q\in\mathcal{Q}}\mathcal{L}(\mathbf{h}^{*},Q_{q,\mathbf{A}^{*}},\mathbf{F}_{T}^{*})-\mathcal{L}(\mathbf{h}^{*},Q^{*},\mathbf{F}_{T}^{*})]$
$\leq$ $2\mathbb{E}[\sup_{(\mathbf{F},Q)\in\mathcal{B}_{r}(B)\circ\mathcal{Q}_{dd^{*}}}|\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\ell(\mathbf{F}\hat{\mathbf{h}}(\mathbf{X}_{T,i})+Q(\mathbf{X}_{T,i})),Y_{T,i})|]+L_{0}\inf_{q\in\mathcal{Q}}\|Q_{q,\mathbf{A}^{*}}-Q^{*}\|_{\infty}$
$\leq$ $2L_{0}\mathbb{E}[\sup_{(\mathbf{F},\mathbf{q},\mathbf{A})\in\mathcal{B}_{r}(B)\circ\mathcal{Q}\circ\mathcal{\mathbf{A}}}|\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}(\mathbf{F}\hat{\mathbf{h}}(\mathbf{X}_{T,i})+\mathbf{q}(\mathbf{A}\mathbf{X}_{T,i})))|]+\frac{2B_{3}}{\sqrt{m}}$
$+L_{0}\inf_{\mathbf{q}\in\mathcal{Q}}\|\mathbf{q}(\mathbf{A}^{*}\cdot)-\mathbf{q}^{*}(\mathbf{A}^{*}\cdot)\|_{\infty}$
[…]
Add $\mathcal{L}_{4}$ and $\mathcal{L}_{5}$ together, we get the wanted result.
$\mathbb{E}[\mathcal{L}(\hat{\mathbf{h}},\hat{Q},\hat{\mathbf{F}}_{T})-\mathcal{L}(\mathbf{h}^{*},Q^{*},\mathbf{F}_{T}^{*})]\precsim\tilde{O}\left(n^{-\frac{\beta}{2(d+1+\beta)}}\mathbb{I}_{\beta\leq 2}+n^{-\frac{\beta}{2d+3\beta}}\mathbb{I}_{\beta>2}\right)+\mathcal{O}\left(m^{-\frac{\beta}{2(d^{*}+1+2\beta)}}\right).$
Usually, by directly assume the transferable condition we can bounding the risk from learning a transferable representation. In our case, it is the term $\mathcal{L}_{4}$. And the term left, which is $\mathcal{L}_{5}$, represents the difficulty of downstream task.
In the condition “complete transfer” , the proof before can process almost the same but get the result $\mathcal{L}_{5}\precsim\frac{1}{\sqrt{m}}$.
Also, when the downstream satisfy the model (8) : $Y_{T}=f^{*}_{T}(\mathbf{h}^{*}(X^{T}))+\varepsilon_{2}$. And by assuming $f^{*}_{T}\in\mathcal{H}^{\beta}$, we can get the total excess risk of downstream task upper bound $\tilde{O}\left(n^{-\frac{\beta}{2(d+1+\beta)}}\mathbb{I}_{\beta\leq 2}+n^{-\frac{\beta}{2d+3\beta}}\mathbb{I}_{\beta>2}\right)+\tilde{O}\left(m^{-\frac{\beta}{2(r+1+2\beta)}}\right)$.
Lemma 9.
$\mathcal{L}_{4}=$ $\mathbb{E}[\mathcal{L}(\hat{\mathbf{h}},Q,\mathbf{F}_{T}^{*})-\mathcal{L}(\mathbf{h}^{*},Q,\mathbf{F}_{T}^{*})]$ (27)
[…]
proof of 9.
We can utilize the property of Wasserstein distance by the distribution of $\mathbf{h}^{*}(\mathbf{X})$ is exactly the same as $\mathcal{U}_{r}$ and A(8) in the Assumption 3 to get
$\mathbb{E}[\mathcal{W}_{1}(\hat{\mathbf{h}}(X_{T}),\mathbf{h}^{*}(X_{T}))]\leq$ $\mathbb{E}[\mathcal{W}_{1}(\hat{\mathbf{h}}(X_{T}),\hat{\mathbf{h}}(X))+\mathcal{W}_{1}(\hat{\mathbf{h}}(X),\mathbf{h}^{*}(\mathbf{X}))+\mathcal{W}_{1}(\mathbf{h}^{*}(\mathbf{X}),\mathbf{h}^{*}(X_{T}))]$
$=$ $\mathbb{E}[\mathcal{W}_{1}(\hat{\mathbf{h}}(X_{T}),\hat{\mathbf{h}}(X))+\mathcal{W}_{1}(\hat{\mathbf{h}}(X),\mathcal{U}_{r})+\mathcal{W}_{1}(\mathbf{h}^{*}(\mathbf{X}),\mathbf{h}^{*}(X_{T}))]$
[…]
The first inequality is by the distance property of Wasserstein distance. The first equality is by the optimality of the function $(f_{0},\mathbf{h}^{*})$. For the second inequality, we can get from lemma 7. Also, the coefficient before $\omega$ is due to the norm constrained of $\hat{\mathbf{h}}$.
Then we denote function $g_{y_{0}}(z)=\ell(z,y_{0})$. By A(7) in Assumption 3, we can show that $g_{y_{0}}(z)$ is a $L_{0}$-lipschitz function.
$\mathcal{L}_{4}=$ $\mathbb{E}_{\mathcal{D}}[\mathcal{L}(\hat{\mathbf{h}},Q,\mathbf{F}_{T}^{*})-\mathcal{L}(\mathbf{h}^{*},Q,\mathbf{F}_{T}^{*})]$
$=$ $\mathbb{E}_{\mathcal{D}}[\mathbb{E}_{Y\sim Y_{T}}[\mathbb{E}_{Z\sim(\mathbf{F}_{T}^{*}\hat{\mathbf{h}}(X_{T})+q(AX_{T}))|Y}[g_{Y}(Z)|Y]-\mathbb{E}_{Z^{\prime}\sim(\mathbf{F}_{T}^{*}\mathbf{h}^{*}(X_{T})+q(AX_{T}))|Y}[g_{Y}(Z^{\prime})|Y]]]$
$\leq$ $\mathbb{E}_{\mathcal{D}}[\mathbb{E}_{Y_{T}}[L_{0}\mathcal{W}_{1}((\mathbf{F}_{T}^{*}\hat{\mathbf{h}}(X_{T})+q(AX_{T}))|Y_{T},(\mathbf{F}_{T}^{*}\mathbf{h}^{*}(X_{T})+q(AX_{T}))|Y_{T})]]$
$=$ $L_{0}\mathbb{E}_{\mathcal{D}}[\mathbb{E}_{Y_{T}}[\mathcal{W}_{1}((\mathbf{F}_{T}^{*}\hat{\mathbf{h}}(X_{T})+q(AX_{T}))|Y_{T},(\mathbf{F}_{T}^{*}\mathbf{h}^{*}(X_{T})+q(AX_{T}))|Y_{T})]]$
$=$ $L_{0}\mathbb{E}_{\mathcal{D}}[\mathbb{E}_{Y_{T}}[\mathcal{W}_{1}(\mathbf{F}_{T}^{*}\hat{\mathbf{h}}(X_{T})|Y_{T},\mathbf{F}_{T}^{*}\mathbf{h}^{*}(X_{T})|Y_{T})]].$
The first inequality is due to the definition of Wasserstein distance. Then by the A(9) in Assumption 3 we can get
$\mathcal{L}_{4}\leq$ $L_{0}\mathbb{E}_{\mathcal{D}}[\mathbb{E}_{Y_{T}}[\mathcal{W}_{1}(\mathbf{F}_{T}^{*}\hat{\mathbf{h}}(X_{T})|Y_{T},\mathbf{F}_{T}^{*}\mathbf{h}^{*}(X_{T})|Y_{T})]]$
$\leq$ $L_{0}\nu\mathbb{E}_{\mathcal{D}}[\mathcal{W}_{1}(\hat{\mathbf{h}}(X_{T}),\mathbf{h}^{*}(X_{T}))]$
$\leq$ $\tilde{O}\left(n^{-\frac{\beta}{2(d+1+\beta)}}\mathbb{I}_{\beta\leq 2}+n^{-\frac{\beta}{2d+3\beta}}\mathbb{I}_{\beta>2}\right).$
C.2 Proof of the Theorem 4
The proof of Theorem 4 is almost the same as proof in Section B.2.
proof of the Theorem 4 .
Following almost the same procedure in the proof of Theorem 2. We consider the case of ”complete transfer” where $Q^{*}=0$, no extra network and the loss $\ell(\cdot,\cdot)$ is the least square form such that ^$\hat{\mathcal{L}}(\hat{\mathbf{h}},\mathbf{F})=\sum_{i=1}^{m}\ell(\mathbf{F}\hat{\mathbf{h}}(\mathbf{x}_{T,i}),Y_{T,i})/m+\chi\|\mathbf{F}\|_{1}=\sum_{i=1}^{m}(\mathbf{F}\hat{\mathbf{h}}(\mathbf{X}_{T,i})-Y_{T,i})^{2}/m+\chi\|\mathbf{F}\|_{1}$.
In this case the ERM solution is a vector $\hat{\mathbf{F}}_{T}\in\arg\min_{\mathbf{F}\in\mathcal{F}}\hat{\mathcal{L}}(\hat{\mathbf{h}},\mathbf{F})$. By the definition of $\hat{\mathbf{F}}_{T}$, we know that ^$\hat{\mathcal{L}}(\hat{\mathbf{F}}_{T},\hat{\mathbf{h}})\leq\hat{\mathcal{L}}(\mathbf{F}_{T}^{*},\hat{\mathbf{h}})$.
$\sum_{i=1}^{m}\frac{1}{m}(Y_{T,i}-\hat{\mathbf{F}}_{T}\hat{\mathbf{h}}(\mathbf{X}_{T,i}))^{2}+\chi\|\hat{\mathbf{F}}_{T}\|_{1}\leq\sum_{i=1}^{m}\frac{1}{m}(Y_{T,i}-F^{*}_{T}\hat{\mathbf{h}}(\mathbf{X}_{T,i}))^{2}+\chi\|\mathbf{F}^{*}_{T}\|_{1}.$
$\sum_{i=1}^{m}\frac{1}{m}(Y_{T,i}-\hat{\mathbf{F}}_{T}\hat{\mathbf{h}}(\mathbf{X}_{T,i}))^{2}\leq\sum_{i=1}^{m}\frac{1}{m}(Y_{T,i}-F^{*}_{T}\hat{\mathbf{h}}(\mathbf{X}_{T,i}))^{2}+\chi\|\hat{\mathbf{F}}_{T}-\mathbf{F}^{*}_{T}\|_{1}.$ (28)
By the equation
$\sum_{i=1}^{m}\frac{1}{m}(Y_{T,i}-\hat{\mathbf{F}}_{T}\hat{\mathbf{h}}(\mathbf{X}_{T,i}))^{2}$ $=\sum_{i=1}^{m}\frac{1}{m}(Y_{T,i}-\mathbf{F}^{*}_{T}\hat{\mathbf{h}}(\mathbf{X}_{T,i})+\mathbf{F}^{*}_{T}\hat{\mathbf{h}}(\mathbf{X}_{T,i})-\hat{\mathbf{F}}_{T}\hat{\mathbf{h}}(\mathbf{X}_{T,i}))^{2}$ (29)
$=\sum_{i=1}^{m}\frac{1}{m}(Y_{T,i}-\mathbf{F}^{*}_{T}\hat{\mathbf{h}}(\mathbf{X}_{T,i}))^{2}$
$\quad+\sum_{i=1}^{m}\frac{2}{m}(Y_{T,i}-\mathbf{F}^{*}_{T}\hat{\mathbf{h}}(\mathbf{X}_{T,i}))(\mathbf{F}^{*}_{T}\hat{\mathbf{h}}(\mathbf{X}_{T,i})-\hat{\mathbf{F}}_{T}\hat{\mathbf{h}}(\mathbf{X}_{T,i}))$
$\quad+\sum_{i=1}^{m}\frac{1}{m}((\mathbf{F}^{*}_{T}-\hat{\mathbf{F}}_{T})\hat{\mathbf{h}}(\mathbf{X}_{T,i}))^{2},$
after estimating the last two term in right part of upper equation. We can replace the left part in (28) by (29) to get the desired result by selecting proper $\chi$.
Then by lemma 10, we know that
$\quad\mathbb{E}_{\mathcal{D}}[\sum_{i=1}^{m}\frac{2}{m}(Y_{T,i}-\mathbf{F}^{*}_{T}\hat{\mathbf{h}}(\mathbf{X}_{T,i}))(\mathbf{F}^{*}_{T}\hat{\mathbf{h}}(\mathbf{X}_{T,i})-\hat{\mathbf{F}}_{T}\hat{\mathbf{h}}(\mathbf{X}_{T,i}))]$
$=\mathbb{E}_{\mathcal{D}}[\sum_{i=1}^{m}\frac{2}{m}(\mathbf{F}^{*}_{T}\mathbf{h}^{*}(\mathbf{X}_{T,i})-\mathbf{F}^{*}_{T}\hat{\mathbf{h}}(\mathbf{X}_{T,i})+\epsilon_{i})(\mathbf{F}^{*}_{T}-\hat{\mathbf{F}}_{T})\hat{\mathbf{h}}(\mathbf{X}_{T,i})]$
$=\mathbb{E}_{\mathcal{D}}[\sum_{i=1}^{m}\frac{2}{m}(\mathbf{F}^{*}_{T}(\mathbf{h}^{*}(\mathbf{X}_{T,i})-\hat{\mathbf{h}}(\mathbf{X}_{T,i}))+\epsilon_{i})(\mathbf{F}^{*}_{T}-\hat{\mathbf{F}}_{T})\hat{\mathbf{h}}(\mathbf{X}_{T,i})]$
$\leq\mathbb{E}_{\mathcal{D}}[\|\mathbf{F}^{*}_{T}-\hat{\mathbf{F}}_{T}\|_{1}\|\sum_{i=1}^{m}\frac{2}{m}\mathbf{F}^{*}_{T}(\mathbf{h}^{*}(\mathbf{X}_{T,i})-\hat{\mathbf{h}}(\mathbf{X}_{T,i}))\hat{\mathbf{h}}(\mathbf{X}_{T,i})\|_{\infty}$
$\quad+\|\mathbf{F}^{*}_{T}-\hat{\mathbf{F}}_{T}\|_{1}\|\sum_{i=1}^{m}\frac{2}{m}\epsilon_{i}\hat{\mathbf{h}}(\mathbf{X}_{T,i})\|_{\infty}]$
$\overset{w.h.p}{\leq}\|\mathbf{F}^{*}_{T}-\hat{\mathbf{F}}_{T}\|_{1}\left(B(\sqrt{\frac{2\log\frac{2}{\delta}}{m}}+B_{2}\gamma_{1}n^{-t(\beta)})+B_{2}\gamma_{1}n^{-t(\beta)}+\sqrt{\frac{2\log\frac{2}{\delta}}{m}}+\sqrt{\frac{2\log\frac{2r}{\delta}}{m}}\right),$
where $\epsilon_{i}$ here is the i.i.d sample of $\varepsilon_{2}$ and for simplicity we omit the footnote.
By A(10) in Assumption 4, we can get
$\mathbb{E}_{\mathcal{D}}[\frac{1}{m}\sum_{i=1}^{m}((\mathbf{F}_{T}^{*}-\hat{\mathbf{F}}_{T})\hat{\mathbf{h}}(\mathbf{X}_{T,i}))^{2}]=\mathbb{E}_{\mathcal{D}}[\|(\mathbf{F}^{*}_{T}-\hat{\mathbf{F}}_{T})\hat{\mathbf{H}}_{T}\|_{2}^{2}]\geq B_{1}\|\mathbf{F}^{*}_{T}-\hat{\mathbf{F}}_{i}\|_{T}^{2}.$
So the (28) can be further deduced as followed:
$B_{1}\|$ $\mathbf{F}^{*}_{T}-\hat{\mathbf{F}}_{T}\|_{2}^{2}\overset{w.h.p}{\leq}\chi\|\hat{\mathbf{F}}_{T}-\mathbf{F}^{*}_{T}\|_{1}$
$+\|\mathbf{F}^{*}_{T}-\hat{\mathbf{F}}_{T}\|_{1}\left(B(\sqrt{\frac{2\log\frac{2}{\delta}}{m}}+B_{2}\gamma_{1}n^{-t(\beta)})+B_{2}\gamma_{1}n^{-t(\beta)}+\sqrt{\frac{2\log\frac{2}{\delta}}{m}}+\sqrt{\frac{2\log\frac{2r}{\delta}}{m}}\right)$
$B_{1}$ $\|\mathbf{F}^{*}_{T}-\hat{\mathbf{F}}_{T}\|_{2}^{2}$
$\overset{w.h.p}{\leq}$ $\Big{(}B(\sqrt{\frac{2\log\frac{2}{\delta}}{m}}+B_{2}\gamma_{1}n^{-t(\beta)})+B_{2}\gamma_{1}n^{-t(\beta)}+\sqrt{\frac{2\log\frac{2}{\delta}}{m}}+\sqrt{\frac{2\log\frac{2r}{\delta}}{m}}+\chi\Big{)}\|\hat{\mathbf{F}}_{T}-\mathbf{F}^{*}_{T}\|_{1}$
$\overset{w.h.p}{\leq}$ $\sqrt{r}\Big{(}B(\sqrt{\frac{2\log\frac{2}{\delta}}{m}}+B_{2}\gamma_{1}n^{-t(\beta)})+B_{2}\gamma_{1}n^{-t(\beta)}+\sqrt{\frac{2\log\frac{2}{\delta}}{m}}+\sqrt{\frac{2\log\frac{2r}{\delta}}{m}}+\chi\Big{)}\|\hat{\mathbf{F}}_{T}-\mathbf{F}^{*}_{T}\|_{2}$
Thus, we have
$B_{1}\|\mathbf{F}^{*}_{T}-$ $\hat{\mathbf{F}}_{T}\|_{2}$
$\overset{w.h.p}{\leq}$ $\sqrt{r}\Big{(}B(\sqrt{\frac{2\log\frac{2}{\delta}}{m}}+B_{2}\gamma_{1}n^{-t(\beta)})+B_{2}\gamma_{1}n^{-t(\beta)}+\sqrt{\frac{2\log\frac{2}{\delta}}{m}}+\sqrt{\frac{2\log\frac{2r}{\delta}}{m}}+\chi\Big{)}$
By set $\chi=\frac{1}{\sqrt{m}}$, we can get the wanted result which means w.p.t that the sparsity of $\hat{\mathbf{F}}_{T}$ is similar with $\mathbf{F}_{T}^{*}$
Lemma 10.
For any $\delta>0$, $i\in[p]$, with probability at least $1-2\delta$ we get:
$\mathbb{E}_{\mathcal{D}}[\|\sum_{i=1}^{m}\frac{1}{m}\mathbf{F}^{*}_{T}(\mathbf{h}^{*}(x_{T,i})-\hat{\mathbf{h}}(x_{T,i}))\hat{\mathbf{h}}(x_{T,i})\|_{\infty}]\leq$ $B(\sqrt{\frac{2\log\frac{2}{\delta}}{m}}+B_{2}\gamma_{1}n^{-t(\beta)}),$
$\mathbb{E}_{\mathcal{D}}[\|\sum_{i=1}^{m}\frac{1}{m}\epsilon_{i}\hat{\mathbf{h}}(x_{T,i})\|_{\infty}]\leq$ $B_{2}\gamma_{1}n^{-t(\beta)}+\sqrt{\frac{2\log\frac{2}{\delta}}{m}}+\sqrt{\frac{2\log\frac{2r}{\delta}}{m}},$
where $\gamma_{1}$ is a constant big enough.
Proof of Lemma 10.
By A(6) in Assumption 3 we know that,
$\|\sum_{i=1}^{m}\frac{1}{m}\mathbf{F}^{*}_{i}(\mathbf{h}^{*}(\mathbf{X}_{T,i})-\hat{\mathbf{h}}(\mathbf{X}_{T,i}))\hat{\mathbf{h}}(\mathbf{X}_{T,i})\|_{\infty}\leq$ $B\sum_{i=1}^{m}\frac{1}{m}\|\mathbf{h}^{*}(\mathbf{X}_{T,i})-\hat{\mathbf{h}}(\mathbf{X}_{T,i})\|_{1},$
$\|\sum_{i=1}^{m}\frac{1}{m}\epsilon_{i}\hat{\mathbf{h}}(\mathbf{X}_{T,i})\|_{\infty}\leq$ $\sum_{i=1}^{m}\frac{1}{m}\|\epsilon_{i}\|_{\infty}\|\mathbf{h}^{*}(\mathbf{X}_{T,i})-\hat{\mathbf{h}}(\mathbf{X}_{T,i})\|_{1}+\|\sum_{i=1}^{m}\frac{1}{m}\epsilon_{i}\mathbf{h}^{*}(\mathbf{X}_{T,i})\|_{\infty},$
By lemma 9, we know that $\mathbb{E}[\mathcal{W}_{1}(\hat{\mathbf{h}}(X_{T}),\mathbf{h}^{*}(X_{T}))]\leq\tilde{O}\left(n^{-\frac{\beta}{2(d+1+\beta)}}\mathbb{I}_{\beta\leq 2}+n^{-\frac{\beta}{2d+3\beta}}\mathbb{I}_{\beta>2}\right)$. For the simplicity of calculation, we can set a constant $\gamma_{1}>$ 0 big enough that $\mathbb{E}[\mathcal{W}_{1}(\hat{\mathbf{h}}(X_{T}),\mathbf{h}^{*}(X_{T}))]\leq\gamma_{1}n^{-t(\beta)}$.
By A(11) in the Assumption 4, $\mathbb{E}[\|\mathbf{h}^{*}(X_{T}^{{}^{\prime}})-\hat{\mathbf{h}}(X_{T}^{{}^{\prime}})\|_{1}]\leq B_{2}n^{-t(\beta)}$. By the independence of dataset $\mathcal{D}$ and $X_{T}^{{}^{\prime}}$. We can use the Hoeffding’s inequality to get:
$\mathbb{P}\left[\left|\sum_{i=1}^{m}\frac{1}{m}\|\mathbf{h}^{*}(\mathbf{X}_{T,i})-\hat{\mathbf{h}}(\mathbf{X}_{T,i})\|_{1}-\mathbb{E}_{X_{T}^{{}^{\prime}}}[\|\mathbf{h}^{*}(X_{T}^{{}^{\prime}})-\hat{\mathbf{h}}(X_{T}^{{}^{\prime}})\|_{1}]\right|\Big{|}\mathcal{D}\right]\geq\delta_{1}]\leq 2\exp\left(-\frac{\delta_{1}^{2}m}{2}\right).$
Then set $\delta_{1}=\sqrt{\frac{2\log\frac{2}{\delta}}{m}}$, for any draw of dataset $\mathcal{D}$ with probability at least $1-\delta$, we know $\sum_{i=1}^{m}\frac{1}{m}\|\mathbf{h}^{*}(\mathbf{X}_{T,i})-\hat{\mathbf{h}}(\mathbf{X}_{T,i})\|_{1}-\mathbb{E}_{X_{T}^{{}^{\prime}}}[\|\mathbf{h}^{*}(X_{T}^{{}^{\prime}})-\hat{\mathbf{h}}(X_{T}^{{}^{\prime}})\|_{1}]]\leq\delta_{1}$. So
$\mathbb{E}_{\mathcal{D}}[\sum_{i=1}^{m}\frac{1}{m}\|\mathbf{h}^{*}(\mathbf{X}_{T,i})-\hat{\mathbf{h}}(\mathbf{X}_{T,i})\|_{1}-\mathbb{E}_{X_{T}^{{}^{\prime}}}[\|\mathbf{h}^{*}(X_{T}^{{}^{\prime}})-\hat{\mathbf{h}}(X_{T}^{{}^{\prime}})\|_{1}]]\leq$ $\delta_{1},$
$\mathbb{E}_{\mathcal{D}}[\sum_{i=1}^{m}\frac{1}{m}\|\mathbf{h}^{*}(\mathbf{X}_{T,i})-\hat{\mathbf{h}}(\mathbf{X}_{T,i})\|_{1}-\mathbb{E}_{X_{T}^{{}^{\prime}}}[\|\mathbf{h}^{*}(X_{T}^{{}^{\prime}})-\hat{\mathbf{h}}(X_{T}^{{}^{\prime}})\|_{1}]]\leq$ $\sqrt{\frac{2\log\frac{2}{\delta}}{m}}.$
Move the term $\mathbb{E}_{X_{T}^{{}^{\prime}}}[\|\mathbf{h}^{*}(X_{T}^{{}^{\prime}})-\hat{\mathbf{h}}(X_{T}^{{}^{\prime}})\|_{1}]$ to the right side of the equation and use the assumption (A5).
$\mathbb{E}_{\mathcal{D}}[\sum_{i=1}^{m}\frac{1}{m}\|\mathbf{h}^{*}(\mathbf{X}_{T,i})-\hat{\mathbf{h}}(\mathbf{X}_{T,i})\|_{1}]\leq$ $\sqrt{\frac{2\log\frac{2}{\delta}}{m}}+\mathbb{E}[\|\mathbf{h}^{*}(X_{T}^{{}^{\prime}})-\hat{\mathbf{h}}(X_{T}^{{}^{\prime}})\|_{1}],$
$\mathbb{E}_{\mathcal{D}}[\sum_{i=1}^{m}\frac{1}{m}\|\mathbf{h}^{*}(\mathbf{X}_{T,i})-\hat{\mathbf{h}}(\mathbf{X}_{T,i})\|_{1}]\leq$ $\sqrt{\frac{2\log\frac{2}{\delta}}{m}}+B_{2}\mathbb{E}[\mathcal{W}_{1}(\mathbf{h}^{*}(X_{T}^{{}^{\prime}}),\hat{\mathbf{h}}(X_{T}^{{}^{\prime}}))],$
$\mathbb{E}_{\mathcal{D}}[\sum_{i=1}^{m}\frac{1}{m}\|\mathbf{h}^{*}(\mathbf{X}_{T,i})-\hat{\mathbf{h}}(\mathbf{X}_{T,i})\|_{1}]\leq$ $\sqrt{\frac{2\log\frac{2}{\delta}}{m}}+B_{2}\gamma_{1}n^{-t(\beta)}.$
Also,
$\mathbb{P}[\|\sum_{j=1}^{m}\frac{1}{m}\epsilon_{i}\mathbf{h}^{*}(\mathbf{X}_{T,i})\|_{\infty}\geq\delta_{2}]$
$=$ $\mathbb{P}[\|\sum_{j=1}^{m}\frac{1}{m}\epsilon_{i}\mathbf{h}^{*}(\mathbf{X}_{T,i})-\mathbb{E}[\varepsilon_{1}\mathbf{h}^{*}(X_{T})]\|_{\infty}\geq\delta_{2}]$
$=$ $\mathbb{P}[\exists k\in[r],|\sum_{j=1}^{m}\frac{1}{m}\epsilon_{i}h^{*}_{k}(\mathbf{X}_{T,i})|\geq\delta_{2}]$
$=$ $\mathbb{P}[\exists k\in[r],|\sum_{j=1}^{m}\frac{1}{m}\epsilon_{i}h^{*}_{k}(\mathbf{X}_{T,i})-\mathbb{E}[\varepsilon_{1}h^{*}_{k}(X)]|\geq\delta_{2}]$
$\leq$ $2r\exp(-\frac{\delta_{2}^{2}m}{2}).$
Then we can set the $\delta_{2}=\sqrt{\frac{2\log\frac{2r}{\delta}}{m}}$ here, then with probability at least $1-\delta$, we know $\|\sum_{j=1}^{m}\frac{1}{m}\epsilon_{i}\mathbf{h}^{*}(\mathbf{X}_{T,i})\|_{\infty}\leq\sqrt{\frac{2\log\frac{2r}{\delta}}{m}}$.
The again by A(11) in the Assumption 4, with probability at least $1-2\delta$ we get:
$\mathbb{E}_{\mathcal{D}}[\|\sum_{i=1}^{m}\frac{1}{m}\mathbf{F}^{*}_{i}(\mathbf{h}^{*}(\mathbf{X}_{T,i})-\hat{\mathbf{h}}(\mathbf{X}_{T,i}))\hat{\mathbf{h}}(\mathbf{X}_{T,i})\|_{\infty}]$
$\leq$ $B\mathbb{E}_{\mathcal{D}}[\sum_{i=1}^{m}\frac{1}{m}\|\mathbf{h}^{*}(\mathbf{X}_{T,i})-\hat{\mathbf{h}}(\mathbf{X}_{T,i})\|_{1}]\leq B(\sqrt{\frac{2\log\frac{2}{\delta}}{m}}+B_{2}\gamma_{1}n^{-t(\beta)}).$
and following the same way before
$\mathbb{E}_{\mathcal{D}}[\|\sum_{i=1}^{m}\frac{1}{m}\epsilon_{i}\hat{\mathbf{h}}(\mathbf{X}_{T,i})\|_{\infty}]\leq$ $\mathbb{E}_{\mathcal{D}}[\sum_{i=1}^{m}\frac{1}{m}\|\epsilon_{i}\|_{\infty}\|\mathbf{h}^{*}(\mathbf{X}_{T,i})-\hat{\mathbf{h}}(\mathbf{X}_{T,i})\|_{1}+\|\sum_{i=1}^{m}\frac{1}{m}\epsilon_{i}\mathbf{h}^{*}(\mathbf{X}_{T,i})\|_{\infty}]$
$\leq$ $B_{2}\gamma_{1}n^{-t(\beta)}+\sqrt{\frac{2\log\frac{2}{\delta}}{m}}+\sqrt{\frac{2\log\frac{2r}{\delta}}{m}}.$
Appendix D Theorem cited during our proof
Theorem 5 (Theorem 3.2 in [83]).
Let $d\in\mathbb{N}$ and $\alpha=s+\beta>0$, where $s\in\mathbb{N}_{0}$ and $\beta\in(0,1]$.
(1) There exists $c>0$ such that for any $K\geq 1$, any $W\geq cK^{(2d+\alpha)/(2d+2)}$ and $L\geq$ $22\left\lceil\log_{2}(d+s)\right\rceil+$ 2,
$\mathcal{E}\left(\mathcal{H}^{\alpha},\mathcal{NN}(W,L,K)\right)\lesssim K^{-\alpha/(d+1)}.$
(2) If $d>2\alpha$, then for any $W,L\in\mathbb{N},W\geq 2$ and $K\geq 1$,
$\mathcal{E}\left(\mathcal{H}^{\alpha},\mathcal{N}\mathcal{N}(W,L,K)\right)\gtrsim(K\sqrt{L})^{-2\alpha/(d-2\alpha)}.$
Theorem 6 (Lemma 2.3 in [83]).
For any $\bm{x}_{1},\ldots,\bm{x}_{n}\in[-B,B]^{d}$ with $B\geq 1$, let $S:=\left\{\left(\phi\left(\bm{x}_{1}\right),\ldots,\phi\left(\bm{x}_{n}\right)\right):\phi\in\mathcal{SNN}_{d,1}(W,L,K)\right\}\subseteq\mathbb{R}^{n}$, then
$\mathcal{R}_{n}(S)\leq\frac{1}{n}K\sqrt{2(L+2+\log(d+1))}\max_{1\leq j\leq d+1}\sqrt{\sum_{i=1}^{n}x_{i,j}^{2}}\leq\frac{BK\sqrt{2(L+2+\log(d+1))}}{\sqrt{n}},$
where $x_{i,j}$ is the $j$-th coordinate of the vector $\tilde{\bm{x}}_{i}=\left(\bm{x}_{i}^{\top},1\right)^{\top}\in\mathbb{R}^{d+1}$. When $W\geq 2$,
$\mathcal{R}_{n}(S)\geq\frac{K}{2\sqrt{2}n}\max_{1\leq j\leq d+1}\sqrt{\sum_{i=1}^{n}x_{i,j}^{2}}\geq\frac{K}{2\sqrt{2n}}.$
Appendix E Experimental details.
In this section, we briefly highlight some important settings in our experiments. More details can be found in our released code in supplementary material.
Datasets. For the classification datasets. VLCS [57] is a challenging dataset consists of 10,700 images. It has VOC2007 (V), LabelMe (L), Caltech (C) and Sun (S) four domains and 5 classes. PACS [58] is a dataset consisting of 9,900 images, which has 7 classes and 4 domains. TerraIncognita [59] is a geographic dataset consisting of images covering a diverse range of environments and objects. It has 4 domains and 24,700 images. OfficeHome [60] contains images collected from four distinct domains: Art (A), Clipart (C), Product (P), and Real-world (R) and corresponding annotations for classification tasks in office environments . It consists of 15,500 images divided into 65 object categories.
For the regression datasets. The tool wear dataset is from the 2010 PHM Challenge [61]. Only three data points, namely C1, C4, and C6, are labeled in the original dataset. The original dataset comprises sensing signals sampled at 50 kHz, encompassing cutting force signals, vibration signals in the x, y, and z dimensions, and acoustic emission signals during processing. The model employs cutting force signals in three directions as input and the average wear values of the three blades for each tool as output to establish transfer tasks.
Hyperparameter search. For the classification task, we maintained a fixed batch size of 64 and utilized the Adam[84] optimizer with the default setting with the only different in learning rate. Both upstream and downstream training consisted of 50 epochs. The upstream training employed a constant learning rate of 0.00001 with the Adam optimizer. For the downstream task, the learning rate for the fully connected layers was set at 0.1, and using the CosineAnnealingLR scheduler. Also, the learning rate for EfficientNet’s feature extractor, which is the Q network, is 0.00001. The hyperparameter search focused solely on the penalty parameters, with values for $\lambda$ selected from {5, 10, 15}, $\tau$ from {0, 10} and $\mu$ from {0, 10}. The hyperparameter for the independence penalty term in the downstream task, $\kappa$ is from {5, 10, 15}.
For the regression task, we maintained a fixed batch size of 50 and employed the Adam optimizer with settings identical to those used in the classification task, except for a variation in the learning rate. The upstream training consisted of 50 epochs and downstream training consisted of 100 epochs. The upstream training employed a learning rate of 0.2. For the downstream task, the learning rate for the fully connected layers was set at 0.02. Both training processes using the CosineAnnealingLR scheduler. The hyperparameter search focused solely on the penalty parameters, with values for $\lambda$ selected from {50, 100, 150, 200}, $\tau$ from {0, 10} and $\mu$ from {0, 10}. The parameters $\kappa$ for the independence penalty term in the downstream task is from {5, 10, 20}.
The size of our fully connected layer for dimensionality reduction is 2048*1024 in classification task, which means the shared feature dimensions in our experiment is 1024. Due to constraints in computational resources and time, we did not explicitly explore the impact of shared feature dimensions $r$ on downstream task performance. This applies to the parameters $\chi$, $\zeta$, and $d^{*}$ as well.
Results of regression dataset. For the improvement in regression dataset is not significant for every domain, so we just list the result here. The phrase ”c4, c6 -c1” in Table 4 represents using the data from domains c4 and c6 as upstream data, and the data from domain c1 as downstream data. The subsequent ”c1, c6 -c4” and ”c1, c4 -c6” have similar meanings.
Table 4: Fine-tuning accuracy on regression dataset and their domains
Transfer task/Algorithm
c4,c6 -c1 MAE RMSE R2
ERM 9.45 13.04 0.71
Ours 3.26 3.91 0.97
c1,c6 -c4 MAE RMSE R2
ERM 4.70 5.68 0.96
Ours 4.65 5.56 0.96
c1,c4 -c6 MAE RMSE R2
ERM 4.31 5.26 0.98
Ours 3.48 4.44 0.98
Details about comparison with other works Due to time constraints and issues with the code logic, we aligned all training details as closely as possible except for the data input method. In the DomainBed codebase[67], the input data is structured such that each remaining domain inputs one batch size of samples, while our input method shuffles all remaining domains and inputs one batch size of samples. Therefore, to align as closely as possible, we set the batch size for upstream training to 21 and the batch size for downstream training to 64 when conducting comparative experiments in the DomainBed codebase. The typical hyperparameters for the compared algorithms are set to their default values in the DomainBed codebase.
Error bar. Due to limited computational resources, we conducted downstream training three times using the checkpoints from the initial upstream training. For comparative results with other methods, particularly those using the DomainBed code [67], we repeated the tests by only varying the seed. The error bar represents the standard deviation of the best result from these experimental runs.
Experiments Compute Resources. Our experiments were conducted on Nvidia DGX Station workstation using a single Tesla V100 GPU unit. For a specific domain in the image dataset, the experimental time for a specific parameter ranges from 0.5 hours to 1.5 hours. Therefore, estimating the experimental results with error bars would require at least three days.
For additional details regarding the small hand-designed model architecture, please refer to the code provided in the supplementary materials.
Appendix F Limitations.
In all the theorems presented in this paper, we analyze the properties of the optimal empirical solutions obtained from our training procedure. However, in practice, it is challenging to ensure that training achieves the optimal empirical solution for most real-world problems solved using neural networks, and this is not unique to our method. Furthermore, in the celebrated paper [85], they give a result about wide (over-parameterized) neural networks satisfy the $PL^{*}$ condition, which explains the (S)GD convergence to a global minimum. Given the large number of parameters in our neural network model (ResNet50 with 25.6M parameters) compared to the sample size ($<$30k), this also implies that our training approach is somehow feasible.